### v1

# Leveraging Temporal Contextualization for Video Action Recognition

Abstract
Pretrained vision-language models have shown effectiveness in video understanding. However, recent studies have not sufficiently leveraged essential temporal information from videos, simply averaging frame-wise representations or referencing consecutive frames. We introduce Temporally Contextualized CLIP (TC-CLIP), a pioneering framework for video understanding that effectively and efficiently leverages comprehensive video information. We propose Temporal Contextualization (TC), a novel layer-wise temporal information infusion mechanism for video that extracts core information from each frame, interconnects relevant information across the video to summarize into context tokens, and ultimately leverages the context tokens during the feature encoding process. Furthermore, our Video-conditional Prompting (VP) module manufactures context tokens to generate informative prompts in text modality. We conduct extensive experiments in zero-shot, few-shot, base-to-novel, and fully-supervised action recognition to validate the superiority of our TC-CLIP. Ablation studies for TC and VP guarantee our design choices. Code is available at https://github.com/naver-ai/tc-clip.
footnotetext: Work done during an internship at NAVER AI Lab.
Corresponding authors.
1 Introduction
Figure 1: Temporal information learning methods. Prior works considered temporal cues during the encoding process via (a) cross-frame attention [22, 32] with [CLS] token interactions or (b) temporal window expansion [33] by adding adjacent frame tokens to key-value pairs. However, the former lacks patch-level interactions, while the latter limits the range of temporal locality. (c) joint space-time attention allows full interactions across all tokens, but it is costly and incurs suboptimal performance (see Fig. 3). (d) Unlike prior approaches, our method aggregates pivotal tokens from a broader range yet efficiently for enhanced temporal integration into key-value pairs.
Pretrained large-scale Vision-Language Models (VLMs) have shown remarkable generalization capability in video understanding and have emerged as promising tools even for zero-shot or open-vocabulary recognition tasks [24, 10, 36]. However, pretraining task-specific models using video-text pairs pose significant challenges, primarily due to substantial computational costs as well as excessive expense for annotated video-text data [31, 34]. Consequently, recent studies in video understanding [30, 22, 13, 32, 26, 33, 5, 9] have shifted their focus towards employing image-based VLMs such as CLIP [24] via fine-tuning—aligning video representations with text embeddings derived from category names.
Despite the successful generalization of CLIP in video recognition, existing approaches fail to fully exploit temporal information in the video feature learning process. Specifically, the token interactions remain restrictive since they either rely on simple averaging of frame-wise representations [26], or attempt to gather information in the temporal axis only through class tokens [32, 22] as illustrated in Fig. 1(a). Although VCLIP [33] incorporates patch-level details by bringing keys and values from neighborhood frames in its self-attention operation, as in Fig. 1(b), its temporal scope is still limited to a local window. Such naïve computation for video representations makes models biased towards static information in their representation learning (e.g., objects and backgrounds) and hampers learning temporal information (e.g., motion and temporal variations). To ensure the global interactions of patch tokens in a spatio-temporal domain, one possible option is to consider every patch token from all frames as a reference during the encoding process as illustrated in Fig. 1(c).
Unfortunately, such a straightforward extension for temporally global interactions in VLMs pretrained with short image-text pairs witnesses extrapolation challenges [23, 4] due to the huge number of tokens in each video. We have observed that a naïve extension of sequence length along the temporal axis degrades its discriminability substantially, as shown in Fig. 2. The joint space-time attention model distributes its attention broadly across the patches but fails to concentrate on informative tokens to recognize actions, incurring suboptimal performance compared to the frame-wise attention baseline. Moreover, this approach suffers from heavy computational overhead due to numerous redundant and similar tokens often corresponding to background regions.
This paper presents Temporally Contextualized CLIP (TC-CLIP), a novel paradigm for extending CLIP to the video domain by encoding holistic video information through advanced temporal analysis. Specifically, our Temporal Contextualization (TC) pipeline summarizes global-range action cues into a small set of tokens for reference during the encoding process. Based on the simple idea by utilizing [CLS] query for self-attention, we could extract informative tokens from each frame and group them contextually. These context tokens act as additional key-value pairs for attention operations in the subsequent layer, presumably serving as a temporal bridge that conveys the video-level context throughout the encoding process. Our preliminary study, shown in Fig. 3, implies that existing methods shown in Fig. 1 offer minimal improvement over the frame-wise attention, highlighting the need for enhanced token interactions.
Figure 2: Pitfall of joint space-time attention. Extending CLIP’s temporal sequence length degrades attention quality, presumably because it was not trained on such long sequences.
Figure 3: Quantitative analysis on temporal modelings. We compare the action recognition performance in the few-shot settings on diverse datasets. All existing methods fall behind our method.
Furthermore, we propose the Video-conditional Prompting (VP) module for the generation of instance-level textual prompts based on produced context tokens. The VP module is composed of cross-attention operations that adopt learnable text prompts as queries and visual temporal tokens as keys and values for injecting video instance representations into video-conditional textual prompts. This strategy alleviates the lack of textual semantics in video action recognition datasets, where category names are the only descriptions of actions (e.g., skateboarding, skydiving, ski jumping), without detailed narratives about acting subjects, observed objects and environment, etc.
[…]
2 Related Work
CLIP for video action recognition. Recent video understanding studies [30, 13, 22, 32, 26, 33, 20, 5, 9] focus on adapting language-image pre-trained models [24, 10, 36]. As a pioneering work, ActionCLIP [30] introduces CLIP [24] into video recognition by adding temporal attention layers after the backbone in a late-fusion fashion. To consider global-range temporal cues during feature encoding, X-CLIP [22] and Vita-CLIP [32] perform cross-frame attention, promoting [CLS] token interaction before spatial attention at each layer. On the other hand, there are lines of work that fully fine-tune CLIP with minimal architectural modifications [26, 33, 5, 9]. ViFi-CLIP [26], a simple fine-tuning method with CLIP without additional temporal modules, generalizes well in open-vocabulary learning. Open-VCLIP [33] expands the temporal view of self-attention by adding keys and values of neighboring frames. Despite the generalization capacity of recent models, we argue they did not adequately leverage temporal information during the video feature encoding process, as they either rely on averaging frame-wise representations [26], [CLS] token interactions [22, 32], or considering only consecutive frames [33]. Unlike prior approaches, we explore a temporal information infusion method to ensure spatio-temporal representation modeling in a broader range.
Token aggregation. Conventional token aggregation methods, prioritizing efficiency enhancement, categorize into hard and soft clusterings. In hard clustering methods like K-Means [21], K-Medoids [15], and DPC-KNN [12], clusters are distinct without overlap, and every data point is assigned solely to one cluster. Bipartite matching [14] divides the data points into two groups and merges a specific number of edges with the highest scores. Soft clustering allows each data point to belong to multiple clusters with varying likelihood. Fuzzy C-Means [7] extends K-Means by enabling non-exclusive partitions and prototypes. Recent token aggregation methods for efficient ViTs [19, 1, 25, 35, 29] leverage aggregation algorithms to reduce the number of tokens for efficiency. While previous approaches focus on efficiency, we employ token aggregation approaches to link relevant tokens across the videos for contextual information.
Prompt learning. Many works have proposed to transfer VLMs to downstream tasks by optimizing a discrete set of prompt vectors [11, 38, 37, 17, 26, 32, 13]. In video recognition, [13] has introduced text prompt tuning, while ViFi-CLIP [26] and Vita-CLIP [32] perform prompting in both vision and text branches. However, these prompt vectors are individually optimized and not shared across the modalities. In image recognition, Co-CoOp [37] performs an instance-conditional prompt tuning by explicitly conditioning the text prompts on [CLS] tokens from image instances. MaPLe [17] learns multi-modal prompting by sharing layerwise context prompts for both branches. Unlikely, we generate video-conditional prompts by utilizing contextualized tokens as vision inputs and injecting summarized video information into text prompt vectors.
3 Proposed Method
3.1 Preliminary
We first review how CLIP [24] is used for video action recognition. In particular, we discuss the encoding procedure based on the vision and text encoders of CLIP, denoted by $\{f_{\theta_{v}},f_{\theta_{c}}\}$, to obtain video and text features, $\{\mathbf{v},\mathbf{c}\}$.
3.1.1 Video encoding.
Given a video $V\in\mathbb{R}^{T\times H\times W\times 3}$ of spatial resolution $H\times W$ with $T$ sampled frames, following the Vision Transformer (ViT) architecture [6], we first divide each frame into $P\times P$ non-overlapping patches and flattening them as a set of vectors $\{\mathbf{x}_{t,i}\in\mathbb{R}^{3P^{2}}\}_{i=1}^{N}$, where $t$ is the frame index, $i$ is the patch index, and $N=HW/P^{2}$ is the number of patches. A sequence of tokens in the $t^{\text{th}}$ frame, denoted by $\mathbf{z}_{t}^{0}$, is fed into the CLIP vision encoder after combining with the spatial positional encoding $\mathbf{e}_{\text{pos}}$ as
$\mathbf{z}_{t}^{0}=[\mathbf{x}_{\text{cls}},\mathbf{x}_{t,1}\mathbf{W}_{\text{emb}},\mathbf{x}_{t,2}\mathbf{W}_{\text{emb}},\ldots,\mathbf{x}_{t,N}\mathbf{W}_{\text{emb}}]+\mathbf{e}_{\text{pos}},$ (1)
where $\mathbf{x}_{\text{cls}}\in\mathbb{R}^{d_{v}}$ is a learnable classification embedding, [CLS] token, and $\mathbf{W}_{\text{emb}}\in\mathbb{R}^{3P^{2}\times d_{v}}$ is a linear projection matrix. The CLIP vision encoder, $f_{\theta_{v}}(\cdot)$, sequentially encodes the frame-level representation $\mathbf{z}_{t}^{l}$ at each layer $l\in\{1,\ldots,L_{v}\}$ as follows:
$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1}),$ (2)
where $f_{\theta_{v}}^{l}(\cdot)$ is the $l^{\text{th}}$ layer of the CLIP vision encoder. Then, we project the [CLS] token of the $t^{\text{th}}$ frame, $\mathbf{z}_{t,0}^{L_{v}}$, onto a common vision-language latent space:
$\mathbf{v}_{t}=\mathbf{z}_{t,0}^{L_{v}}\mathbf{W}_{\text{vis}},$ (3)
where $\mathbf{W}_{\text{vis}}\in\mathbb{R}^{d_{v}\times d_{vl}}$ is a projection matrix. Finally, the video representation $\mathbf{v}$ is obtained by averaging the per-frame representations $\mathbf{v}_{t}$ as
$\mathbf{v}=\text{AvgPool}([\mathbf{v}_{1},...,\mathbf{v}_{T}]).$ (4)
3.1.2 Text encoding.
Given a text description $C$ for a video, the input of the text encoder, $\textbf{c}^{0}$, is obtained by tokenizing words in the description and computing their word embedding vectors. In addition to the embeddings from category names, one can augment a sequence of prompt embeddings $\textbf{p}^{0}$, which are obtained from either hand-crafted templates such as “a photo of a” or learnable prompt vectors. The CLIP text encoder, $f_{\theta_{c}}(\cdot)$, sequentially processes a sequence of text embeddings including prompt embeddings, which is denoted by $[\textbf{p}^{0},\textbf{c}^{0}]$, and computes an intermediate representation at each layer $l\in\{1,\ldots,L_{c}\}$ as follows:
$[\mathbf{p}^{l},\mathbf{c}^{l}]=f_{\theta_{c}}^{l}([\mathbf{p}^{l-1},\mathbf{c}^{l-1}]),$ (5)
where $f_{\theta_{c}}^{l}(\cdot)$ denotes the $l^{\text{th}}$ layer of the CLIP text encoder. The final text representations $\mathbf{c}$ is obtained by projecting the [EOS] token from the last layer to the vision-language latent space using a matrix $\mathbf{W}_{\text{text}}\in\mathbb{R}^{d_{l}\times d_{vl}}$, i.e., $\mathbf{c}=\mathbf{c}_{\text{eos}}^{L_{c}}\mathbf{W}_{\text{text}}$.
3.1.3 Video-text alignment.
Finally, the similarity between video and text embeddings are formulated as follows:
$\mathrm{sim}(\mathbf{v},\mathbf{c})=\frac{\langle\mathbf{v},{\mathbf{c}\rangle}}{\lVert\mathbf{v}\rVert\lVert\mathbf{c}\rVert}.$ (6)
During training, the goal is to maximize the similarity if $V$ and $C$ are matched but minimize it otherwise. For inference, the category with the highest similarity is chosen as the prediction output.
3.2 Temporal Contextualization (TC)
To alleviate the chronic drawbacks of existing CLIP-based video representations and leverage the spatio-temporal context in an entire video during the encoding process, we propose Temporal Contextualization (TC). TC consists of three steps, which are 1) informative token selection in each frame, 2) context summarization across spatio-temporal dimensions, and 3) temporal context infusion to all tokens in the subsequent layer. Fig. 4 illustrates an overview of TC. We discuss the three steps of our TC pipeline in the following.
Figure 4: Overview of Temporal Contextualization (TC). (a) Starting with the original patch tokens of the entire video, we first select informative seed tokens from each frame and perform temporal context summarization to aggregate relevant tokens to obtain context tokens. (b) The encoded context tokens are subsequently utilized as key-value pairs for self-attention in the next layer. (c) We employ attention score with [CLS] tokens as the criterion for seed token selection.
3.2.1 Informative token selection.
Due to a large number of redundant patches, the use of all patches for encoding a video may be suboptimal for extracting desired temporal information. Our approach focuses on selecting the informative tokens before performing the token aggregation. To this end, the attention scores obtained from self-attention operations in each frame are used as criteria to select seed tokens. Specifically, given patch tokens $\{\mathbf{z}_{t,i}\}_{i=1}^{N}$ in the $t^{\text{th}}$ frame, a set of attention scores $\{\mathbf{a}(\mathbf{z}_{t,i})\}_{i=1}^{N}$ is driven from the attention of the query of the [CLS] token to the keys of patch tokens, which is given by
$\mathbf{a}(\mathbf{z}_{t})=\text{Softmax}\Big{(}\frac{\mathbf{q}_{\text{cls}}\mathbf{K}_{\mathbf{z}_{t}}^{\mathsf{T}}}{\sqrt{d}}\Big{)},$ (7)
where both the query $\mathbf{q}_{\text{cls}}=\mathbf{z}_{\text{t,0}}\mathbf{W}_{q}\in\mathbb{R}^{d}$ and keys $\mathbf{K}_{\mathbf{z}_{t}}=\mathbf{z}_{t}\mathbf{W}_{k}\in\mathbb{R}^{(N+1)\times d}$ are given by linear projections of input $\mathbf{z}_{t}\in\mathbb{R}^{(N+1)\times d}$. In practice, our model yields multiple [CLS] attention scores through multi-head self-attention (MHSA) operations and compute the average of the attention scores from all heads, i.e., $\bar{\mathbf{a}}_{t,i}=\sum_{h=1}^{H}\mathbf{a}_{t,i}^{h}/H$, where $\mathbf{a}_{t,i}^{h}=\mathbf{a}^{h}(\mathbf{z}_{t,i})$ is the attention score for the $i^{\text{th}}$ patch $\mathbf{z}_{t,i}$ in the $t^{\text{th}}$ frame and $H$ is a number of heads. Finally, we formulate a set of seed tokens for the $t^{\text{th}}$ frame, $\mathcal{S}_{t}$, by selecting the indices for $n_{s}$ elements with highest attention score $\bar{\mathbf{a}}_{t,i}$, where the number of seed tokens in each frame $n_{s}$ is controlled by the ratio of $\alpha=n_{s}/N$.
3.2.2 Temporal context summarization.
We now interconnect the seed tokens derived from individual frames based on their relevance and identify a collection of context tokens. Specifically, we collect all the seed tokens from all frames as $\mathcal{S}=\{(t,i)|i\in\mathcal{S}_{t},t=1,\ldots,T\}$ and perform their spatio-temporal summarization by clustering and merging all the seed tokens in $\mathcal{S}$ as follows:
$\hat{\mathbf{s}}=\phi\big{(}\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}\big{)}.$ (8)
where $\hat{\mathbf{s}}\in\mathbb{R}^{k\times D}$ denotes a collection of the summarized tokens, which we call context tokens, $\phi$ is a token aggregation function, and $\hat{\textbf{z}}_{t,i}$ indicates an interim token encoded from $\textbf{z}_{t,i}$ via the self-attention operation. In practical, we construct a bipartite graph whose vertices and edges represent the original seed tokens and their pairwise similarities. Then, we iteratively perform bipartite matching [14] that splits the vertex set and merges $r$ pairs of vertices based on their similarities by simple averaging of vertex features until there remain $k$ number of tokens. Subsequently, the context tokens $\hat{\mathbf{s}}$ are fed into a feed-forward network (FFN) as done in the original encoding procedure and are forwarded to the next layer.
3.2.3 Temporal context infusion.
Finally, we infuse the summarized context to all patch tokens by modifying the self-attention function. The keys and values of self-attention in every frame are expanded to include context tokens as follows:
$\text{Attention}_{\text{TC}}(\mathbf{z}_{t},\mathbf{s})$ $=\text{Softmax}\Big{(}\frac{\mathbf{Q}_{\mathbf{z}_{t}}\big{[}\mathbf{K}_{\mathbf{z}_{t}}|\mathbf{K}_{\mathbf{s}}\big{]}^{\mathsf{T}}}{\sqrt{d}}+\mathbf{B}\Big{)}\big{[}\mathbf{V}_{\mathbf{z}_{t}}|\mathbf{V}_{\mathbf{s}}\big{]},$ (9)
where $\mathbf{K}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{k}$ and $\mathbf{V}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{v}$ are linear projections of the context tokens $\mathbf{s}\in\mathbb{R}^{k\times d}$. Here, $\mathbf{B}\in\mathbb{R}^{(N+1)\times(N+k+1)}$ is a bias matrix that distinguishes between frame-level local information and video-level global information in the expanded key matrix as
$\mathbf{B}_{ij}=\begin{cases}b_{\text{local}}&\text{if }j\leq N+1\\ b_{\text{global}}&\text{otherwise},\end{cases}$ (10)
where $b_{\text{local}}$ and $b_{\text{global}}$ are learnable parameters and defined for multiple heads at each layer. We build our TC pipeline in a layer-wise manner, and thus the encoding process of each layer is expressed as
$\hat{\mathbf{z}}_{t}^{l}$ =$=\begin{cases}\text{MHSA}(\text{LN}(\mathbf{z}_{t}^{l-1}))+\mathbf{z}_{t}^{l-1}&\text{if }l=1\\ \text{MHSA}_{\text{TC}}(\text{LN}(\mathbf{z}_{t}^{l-1}),\text{LN}(\mathbf{s}^{l-1}))+\mathbf{z}_{t}^{l-1}&\text{otherwise},\end{cases}$ (11)
$\mathbf{z}_{t}^{l}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{z}}_{t}^{l}))+\hat{\mathbf{z}}_{t}^{l},$ (12)
$\mathbf{s}^{l}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{s}}^{l}))+\hat{\mathbf{s}}^{l},$ (13)
where $\text{MHSA}_{\text{TC}}(\cdot,\cdot)$ denotes the TC-based MHSA operation including Eq. (9) and $\text{LN}(\cdot)$ stands for the layer normalization function.
3.3 Video-conditional Prompting (VP)
We present the Video-conditional Prompting (VP) module, which aims to further leverage the precisely aggregated temporal information in visual domain for text encoding. We impose a cross-attention between prompt vectors and context tokens to enrich the information in the prompt vectors, and the detailed procedure is illustrated in Fig. 5. Let $\textbf{c}^{l-1}$ and $\textbf{p}^{l-1}$ be class name tokens and learnable prompt vectors from the $(l-1)^{\text{th}}$ layer of the text encoder, respectively. We derive temporally contextualized prompt vectors $\hat{\mathbf{p}}^{l-1}$ by passing the layer-normalized prompt tokens and context tokens through a cross-attention layer as follows:
Figure 5: Video-conditional Prompting (VP) module. Video information from the context tokens is injected to the text prompt vectors based on a cross-attention mechanism, generating instance-level prompts that support the lack of textual semantics.
$\mathbf{s}_{\text{proj}}^{l}$ $=\text{SG}(\mathbf{s}^{l}\mathbf{W}_{\text{vis}}),$ (14)
$\hat{\mathbf{p}}^{l-1}$ $=\text{MHCA}(\text{LN}_{p}(\mathbf{p}^{l-1}),\text{LN}_{s}(\mathbf{s}_{\text{proj}}^{l}))+\mathbf{p}^{l-1}$ (15)
where $\text{SG}(\cdot)$ is a stop-gradient function, $\mathbf{W}_{\text{vis}}$ is a weight matrix of CLIP to linearly project vision representations to a common vision-language latent space, and $\text{MHCA}(\cdot,\cdot)$ is a multi-head cross-attention opearation for interactions across modalities, accepting text prompt vectors as queries and vision features as keys and values. Finally, the contextualized prompt vector $\hat{\mathbf{p}}^{l-1}$ are updated by FFN:
$\tilde{\mathbf{p}}^{l-1}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{p}}^{l-1})+\hat{\mathbf{p}}^{l-1}$ (16)
The VP module $f_{\theta_{\text{VP}}}(\cdot,\cdot)$ is defined by a composition of Eq. (15) and Eq. (16), and executed before the last layer of text encoder $f_{\theta_{c}}$. Finally, the new formulation of our encoding process in the text modality is given by
$[\mathbf{p}^{l},\mathbf{c}^{l}]=\begin{cases}f_{\theta_{c}}^{l}([f_{\theta_{\text{VP}}}(\mathbf{p}^{l-1},\mathbf{s}_{\text{proj}}^{l}),\mathbf{c}^{l-1}])&\text{if}~{}l=L_{c}\\ f_{\theta_{c}}^{l}([\mathbf{p}^{l-1},\mathbf{c}^{l-1}])&\text{otherwise.}\end{cases}$ (17)
The resulting instance-conditioned prompt vectors effectively support the lack of textual semantics in action recognition datasets. The impact of our VP design is verified in the experiments in Sec. 5.
3.4 Training Objective
TC-CLIP learns to maximize the similarity of video representations $\mathbf{v}$ and text representations $\mathbf{c}$ for matching pairs in a mini-batch via the cross-entropy loss:
$\mathcal{L}=-\sum_{i}\log\frac{\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{j})/\tau)},$ (18)
where $\tau$ is a learnable temperature parameter. Our model is fully fine-tuned in an end-to-end manner.
4 Experiments
4.1 Experimental Setup
Datasets. We conduct experiments over 5 video benchmarks: Kinetics-400 [16] & 600 [2], HMDB-51 [18], UCF-101 [28], and Something-Something v2 (SSv2) [8]. Following [26], our evaluation protocol includes zero-shot, few-shot, base-to-novel generalization, and fully-supervised action recognition.
Implementation details. We adopt CLIP with ViT-B/16 for all experiments. Our baseline is ViFi-CLIP [26]. During the training, we sample 16 frames to form a video clip. During the evaluation, two temporal clips with one spatial crop (2 $\times$ 1 view) per video are sampled to produce a prediction unless otherwise stated. For training recipes, we follow [26] for zero-shot, few-shot, and fully-supervised settings and follow [9] for base-to-novel generalization. Following [33, 20, 5, 9], we adopt linear weight-space ensembling of our fine-tuned model and CLIP for the zero-shot protocol. We set the number of merged tokens $r$ in each bipartite matching iteration to 100, and the weight of the VP module is initialized with the weight from the last layer of the CLIP text encoder. All models are trained using 4 NVIDIA Tesla V100 GPUs. See supplementary material for more details.
Table 1: Comparison with state-of-the-arts on zero-shot action recognition. All the models are trained on Kinetics-400 and directly evaluated on other datasets. † denotes the methods with weight-space ensemble. Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
Method HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
Vanilla CLIP [24] 40.8 $\pm$ 0.3 63.2 $\pm$ 0.2 59.8 $\pm$ 0.3 83.5 $\pm$ 0.2 54.6
ActionCLIP [30] 40.8 $\pm$ 5.4 58.3 $\pm$ 3.4 66.7 $\pm$ 1.1 91.6 $\pm$ 0.3 55.3
A5 [13] 44.3 $\pm$ 2.2 69.3 $\pm$ 4.2 55.8 $\pm$ 0.7 81.4 $\pm$ 0.3 56.5
X-CLIP [22] 44.6 $\pm$ 5.2 72.0 $\pm$ 2.3 65.2 $\pm$ 0.4 86.1 $\pm$ 0.8 60.6
Vita-CLIP [32] 48.6 $\pm$ 0.6 75.0 $\pm$ 0.6 67.4 $\pm$ 0.5 - 63.7
ViFi-CLIP [26] 51.3 $\pm$ 0.6 76.8 $\pm$ 0.7 71.2 $\pm$ 1.0 92.2 $\pm$ 0.3 66.4
Open-VCLIP [33] † 53.9 $\pm$ 1.2 83.4 $\pm$ 1.2 73.0 $\pm$ 0.8 93.2 $\pm$ 0.1 70.1
TC-CLIP (Ours) † 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 94.4 $\pm$ 0.2 71.0
Using LLM-based text augmentation
MAXI [20] † 52.3 $\pm$ 0.7 78.2 $\pm$ 0.8 71.5 $\pm$ 0.8 92.5 $\pm$ 0.4 67.3
OST [5] † 55.9 $\pm$ 1.2 79.7 $\pm$ 1.1 75.1 $\pm$ 0.6 94.6 $\pm$ 0.2 70.2
FROSTER [9] † 54.8 $\pm$ 1.3 84.8 $\pm$ 1.1 74.8 $\pm$ 0.9 - 71.5
TC-CLIP (Ours) † 56.0 $\pm$ 0.3 85.4 $\pm$ 0.8 78.1 $\pm$ 1.0 95.7 $\pm$ 0.3 73.2
4.2 Quantitative Comparison
We mainly compare our method with CLIP-based video recognition models: Vanilla CLIP [24], ActionCLIP [30], A5 [13], X-CLIP [22], Vita-CLIP [32], ViFi-CLIP [26], Open-VCLIP [33], OST [5], and FROSTER [9]. To fairly compare our method with approaches that utilize Large Language Model (LLM)-based text augmentation [20, 5, 9], we divide our results into two versions: one using the original action category names (colored with blue) and the other adopting the LLM-rephrased category names from FROSTER [9] (colored with purple). Note that experiments on the SSv2 dataset do not involve LLM-rephrasing.
[…]
Table 2: Comparison with state-of-the-arts on few-shot action recognition. All the models are directly fine-tuned from CLIP. Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
HMDB-51 UCF-101 SSv2 All
Method $K$=$2$ $K$=$4$ $K$=$8$ $K$=$16$ $K$=$2$ $K$=$4$ $K$=$8$ $K$=$16$ $K$=$2$ $K$=$4$ $K$=$8$ $K$=$16$  Avg.
Vanilla CLIP [24] 41.9 41.9 41.9 41.9 63.6 63.6 63.6 63.6 2.7 2.7 2.7 2.7 36.1
ActionCLIP [30] 47.5 57.9 57.3 59.1 70.6 71.5 73.0 91.4 4.1 5.8 8.4 11.1 46.5
A5 [13] 39.7 50.7 56.0 62.4 71.4 79.9 85.7 89.9 4.4 5.1 6.1 9.7 46.8
X-CLIP [22] 53.0 57.3 62.8 64.0 76.4 83.4 88.3 91.4 3.9 4.5 6.8 10.0 50.2
ViFi-CLIP [26] 57.2 62.7 64.5 66.8 80.7 85.1 90.0 92.7 6.2 7.4 8.5 12.4 52.9
TC-CLIP (Ours) 57.3 62.3 67.3 68.6 85.9 89.9 92.5 94.6 7.3 8.6 9.3 14.0 54.8
Using LLM-based text augmentation
OST [5] 59.1 62.9 64.9 68.2 82.5 87.5 91.7 93.9 7.0 7.7 8.9 12.2 53.9
TC-CLIP (Ours) 58.6 63.3 65.5 68.8 86.8 90.1 92.0 94.3 7.3 8.6 9.3 14.0 54.9
Table 3: Comparison with state-of-the-arts on base-to-novel generalization. All the models are directly fine-tuned from CLIP. † denotes the results are taken from [9]. Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
K-400 HMDB-51 UCF-101 SSv2 All (Avg.)
Method Base Novel HM Base Novel HM Base Novel HM Base Novel HM Base Novel HM
Vanilla CLIP [24] 62.3 53.4 57.5 53.3 46.8 49.8 78.5 63.6 70.3 4.9 5.3 5.1 49.8 42.3 45.7
ActionCLIP [30] 61.0 46.2 52.6 69.1 37.3 48.5 90.1 58.1 70.7 13.3 10.1 11.5 58.5 37.9 46.0
A5 [13] 69.7 37.6 48.8 46.2 16.0 23.8 90.5 40.4 55.8 8.3 5.3 6.4 53.7 24.8 33.9
X-CLIP [22] 74.1 56.4 64.0 69.4 45.5 55.0 89.9 58.9 71.2 8.5 6.6 7.4 60.5 41.9 49.5
ViFi-CLIP [26] 76.4 61.1 67.9 73.8 53.3 61.9 92.9 67.7 78.3 16.2 12.1 13.9 64.8 48.6 55.5
Open-VCLIP [33] † 76.5 62.6 68.9 70.3 50.4 58.9 94.8 77.5 85.3 16.0 11.0 13.0 64.4 50.4 56.5
TC-CLIP (Ours) 78.9 63.6 70.4 73.3 54.1 62.2 95.5 78.0 85.9 17.5 13.4 15.2 66.3 52.3 58.5
Using LLM-based text augmentation
FROSTER [9] 77.8 64.3 70.4 74.1 58.0 65.1 95.3 80.0 87.0 18.3 12.2 14.6 66.4 53.6 59.3
TC-CLIP (Ours) 79.1 65.4 71.6 73.3 59.1 65.5 95.4 81.6 88.0 17.5 13.4 15.2 66.3 54.9 60.1
Few-shot action recognition. We demonstrate the learning capacity of our method under a challenging few-shot scenario. In this setting, models are directly fine-tuned from CLIP on HMDB-51, UCF-101, and SSv2 using $K$-shot samples, where $K$ is set to 2, 4, 8, and 16. Table 2 shows the comparison with state-of-the-art video action recognition methods on the few-shot setting. Our method achieves state-of-the-art performance with an average accuracy of 54.8%, surpassing our baseline model ViFi-CLIP [26] by 1.9%p.
Base-to-novel generalization. Similar to the few-shot setting, models are directly fine-tuned from CLIP on the base classes of K-400, HMDB-51, UCF-101, and SSv2, and evaluated for both base and novel classes. The top-1 accuracy on base and novel classes, and their harmonic mean (HM) are reported. Table 3 shows that our model achieves state-of-the-art performance on the novel classes and HM across all datasets. In particular, TC-CLIP achieves an average HM accuracy of 58.5%, surpassing previous state-of-the-arts by 2.0%p.
Table 4: Comparison with state-of-the-arts on fully-supervised action recognition on Kinetics-400. Views are denoted with temporal clips $\times$ spatial crops. † means that computational costs are measured in our environment using a single V100 GPU. The results of other methods are taken from [26].
Method Top-1 Top-5 Frames Views GFLOPs Params (M)
ActionCLIP [30] 83.8 96.2 32 10 $\times$ 3 563 168.5
X-CLIP [22] 84.7 96.8 16 4 $\times$ 3 287 131.5
Vita-CLIP [22] † 82.9 96.3 16 4 $\times$ 3 307 185.3
ViFi-CLIP [26] 83.9 96.3 16 4 $\times$ 3 281 124.7
OST [5] 83.2 - 16 1 $\times$ 1 - -
TC-CLIP (Ours) 85.2 96.9 16 4 $\times$ 3 304 127.5
Fully-supervised action recognition. Table 4 shows performance comparison results under fully-supervised setting, where the models are both trained and evaluated on the K-400 dataset. Our model achieves top-1 accuracy of 85.2% in the K-400 validation split, improving 1.3%p over our baseline ViFi-CLIP [26].
Table 5: Component-wise ablation on the zero-shot setting. ($\Delta$) denotes the average top-1 accuracy improvement over baseline.
Case TC VP WE HMDB-51 UCF-101 K-600 All ($\Delta$)
Baseline 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 67.3
(a) w/o Video-conditional Prompting (VP) 54.3 $\pm$ 0.6 81.9 $\pm$ 1.0 75.5 $\pm$ 0.5 70.6 ($+$3.3)
(b) w/o Temporal Contextualization (TC) 53.2 $\pm$ 0.8 82.0 $\pm$ 0.9 74.7 $\pm$ 0.7 70.0 ($+$2.7)
(c) w/o Weight-space Ensembling (WE) 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 68.9 ($+$1.6)
TC-CLIP 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 71.0 ($+$3.7)
Table 6: Temporal Contextualization (TC) ablation on the few-shot setting. SE, JE, and Bias denote spatial embedding, joint space-time embedding, and bias (Eq. 10). Default settings are marked in gray.
(a) Seed token selection metric.
Case HMDB UCF SSv2 All
No selection 62.8 89.8 9.7 54.1
Head-wise key norm 62.3 89.8 9.8 54.0
Averaged key norm 62.5 89.4 9.3 53.7
Head-wise CLS attention 63.4 89.9 9.7 54.3
Averaged CLS attention 63.4 90.2 9.9 54.5
(b) Token aggregation method.
Case HMDB UCF SSv2 All
No merge 57.2 85.6 7.7 50.2
Temporal tube 57.9 86.4 8.3 50.9
Random merge 58.8 87.1 7.5 51.2
K-means 61.7 89.7 9.0 53.4
Bipartite matching 63.4 90.2 9.9 54.5
(c) Seed token ratio $\alpha$.
$\alpha$ HMDB UCF SSv2 All
0.2 62.6 90.1 9.8 54.2
0.3 63.4 90.2 9.9 54.5
0.4 63.2 90.4 9.8 54.5
0.5 63.3 90.3 9.8 54.5
0.6 63.1 90.2 9.8 54.4
(d) Context token $k$.
$k$ HMDB UCF SSv2 All
16 63.1 89.3 9.1 53.8
32 63.6 89.9 9.4 54.3
64 63.7 90.1 9.7 54.5
96 63.4 90.2 9.9 54.5
128 62.8 90.1 9.9 54.3
256 60.9 89.5 9.3 53.2
(e) Positional embedding.
Case HMDB UCF SSv2 All
SE 62.9 90.0 9.8 54.2
JE 63.2 90.2 9.8 54.4
SE+Bias 63.4 90.2 9.9 54.5
JE+Bias 62.9 90.2 9.8 54.3
5 Analysis and Discussion
This section examines the design choices and impact of each component in our model: Temporal Contextualization (TC) and Video-conditional Prompting (VP). We mainly adopt the zero- and few-shot settings and report the average of top-1 accuracy from $K=2,4,8,16$ for the $K$-shot setup.
5.1 Component-wise Ablation
Table 5 shows the impact of TC, VP, and weight-space ensembling (WE) on our baseline in the zero-shot setting. (a) Adopting TC with WE leads to an average gain of 3.3%p, while (b) incorporating VP with WE results in a 2.7%p improvement, highlighting the significance of TC in our framework. (c) When both VP and TC are applied to the baseline, we observed a 1.6%p improvement on average. Finally, by integrating all three components, TC-CLIP achieves a total 3.7%p gain on average.
5.2 Temporal Contextualization (TC) Ablation
Table 6 analyzes the design choice of TC. Experiments are conducted on the few-shot setting using the baseline model with TC.
Seed token selection metric. Using [CLS] attention score averaged from multiple heads is the best choice, resulting in a 0.4%p gain on average compared to the no-selection case, i.e., using all tokens for aggregation.
Token aggregation methods. Using seed tokens without merging leads to inferior performance, encountering the extrapolation issue. When comparing naive aggregation methods, random merge surpasses temporal tube since random merge has more chance to aggregate relevant tokens while temporal tube are enforced to group tokens of the same spatial positions. Finally, we compared the clustering algorithms and selected bipartite matching as our default choice.
Seed token selection ratio $\alpha$. While TC is not sensitive to the choice of $\alpha$, we picked $\alpha=0.3$ as our default value, i.e., using 30% of total tokens.
Context token number $k$. By varying the value of $k$, we observed that selecting a modest amount of merging degree is important for achieving the best performance. An excessively low $k$ leads to the lost of spatial details like part-wise information, while an overly high $k$ results in inadequate contextualization of tokens into relevant sets. Therefore, we choose $k=96$ as our default setting.
Positional embedding. Among spatial embedding (SE), joint space-time embedding (JE), and our proposed learnable bias (Eq. 10), we observe that using SE with bias yields the best result. We conjecture that the bias effectively consolidates the local frame-level information and global video-level information in a layer-wise and head-wise manner.
Table 7: Video-conditional Prompting (VP) ablation. Default settings are marked in gray.
(a) Vision token selection (few-shot).
Case HMDB UCF SSv2 All
CLS tokens 63.3 90.3 9.8 54.4
GAP tokens 63.7 90.5 9.8 54.7
Context tokens 63.9 90.7 9.8 54.8
(b) Number of prompt vectors (few-shot).
Case HMDB UCF SSv2 All
2 62.0 90.1 9.4 53.8
4 63.9 90.7 9.8 54.8
8 63.7 90.4 9.8 54.7
(c) Text prompting design (zero-shot). All the models are evaluated without weight ensembling.
Case Use context tokens? HMDB-51 UCF-101 K-600 All ($\Delta$)
Baseline 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 67.3
(a) Learnable prompt vectors 52.4 $\pm$ 0.4 78.4 $\pm$ 1.3 70.6 $\pm$ 0.7 67.1 (--0.2)
(b) Video-conditional prompting 53.2 $\pm$ 0.8 80.4 $\pm$ 0.7 71.6 $\pm$ 0.9 68.4 (+1.1)
(c) Video-conditional prompting 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 68.9 (+1.6)
(d) Vision-text late-fusion 53.7 $\pm$ 0.7 79.0 $\pm$ 0.7 70.9 $\pm$ 0.6 67.9 (+0.6)
5.3 Video-conditional Prompting (VP) Ablation
In Table 7, we analyze the design choice of the VP module in TC-CLIP on the few-shot and zero-shot setting.
Vision token selection. We compare using the context tokens with using [CLS] or Global Average Pooled (GAP) tokens from every frame. Using context tokens yields the best result, demonstrating that the proper contextualization from the video is essential for transferring the video information to the text side.
Number of prompt vectors. Increasing the number of prompt vectors does not necessarily improve performance. 4 prompt vectors are employed by default.
Text prompting design. We observe that (a) naively adding learnable prompt vectors without video instance conditioning does not enhance the zero-shot transferability and decreases the average accuracy by 0.2%p. (b) In contrast, employing VP design with [CLS] tokens consistently improves the accuracy across all datasets, and (c) using context tokens further enhances the performance, resulting in a 1.6%p gain. We also compare VP with (d) vision-text late-fusion design, i.e., the cross-attention of context tokens and the final representation of the text embedding. This design shows worse performance in UCF-101 and K-600 datasets than our VP, demonstrating the superiority of our design choice.
5.4 Qualitative Analysis
Context token visualization. Fig. 6 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP. In this video, the informative regions regarding the action of “disc golfing” in each frame, including the person and the disc, are selected as seed tokens. To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region. Note that a single context token (highlighted in red) successfully tracks the disc across multiple frames.
Attention visualization. Fig. 7 visualizes the attention map of ViFi-CLIP [26] and our TC-CLIP on the SSv2 dataset. In this video where two hands grab objects and then move away, ViFi-CLIP [26] fails to attend to the hands from the middle of the sequence and misinterprets the action as “colliding with each other”. In contrast, our TC-CLIP considers the temporal context across the sequence by its design, and thus consistently attends to the hands throughout the entire video and correctly predicts the action as “moving away from each other”. See more visualizations in Appendix 0.E.
Figure 6: Context token visualization. TC-CLIP selects the informative seed tokens and summarizes them into context tokens across frames. The disc (red) is merged into one token over the video.
Figure 7: Attention visualization. While ViFi-CLIP fails to attend to the movement of hands moving farther away and misinterprets the action as colliding, TC-CLIP correctly predicts with temporal consistency.
6 Conclusion
We have introduced TC-CLIP, a novel video understanding paradigm that leverages holistic video information within the encoding process. Unlike prior approaches that access only a limited range of tokens, our temporal contextualization summarizes informative tokens from the entire video and utilizes them during the attention operation. While these tokens are exploited to infuse temporal information on the vision side, they also serve as a source for video-conditional text prompting, enhancing instance-wise context on the text side. Extensive experiments and analysis on diverse video recognition benchmarks and evaluation protocols demonstrate the superiority of TC-CLIP and its design choices.
Limitation. Due to resource constraints, we verified TC-CLIP only using ViT-B. We will further explore its scalability with ViT-L and ViT-H.
Appendix 0.A Fine-tuning with the Kinetics-400 Pretrained Model
Table 8: Comparison with state-of-the-arts on few-shot action recognition using Kinetics-400 pretrained model. All the models are first pretrained on Kinetics-400 and subsequently fine-tuned on each dataset.
HMDB-51 UCF-101 SSv2 All
Method $K$=$2$ $K$=$4$ $K$=$8$ $K$=$16$ $K$=$2$ $K$=$4$ $K$=$8$ $K$=$16$ $K$=$2$ $K$=$4$ $K$=$8$ $K$=$16$  Avg.
ActionCLIP [30] 54.3 56.2 59.3 66.1 76.7 80.4 87.6 91.8 4.8 6.9 9.1 12.3 50.5
A5 [13] 46.7 50.4 61.3 65.8 76.3 84.4 90.7 93.0 4.5 6.7 7.2 9.5 49.7
X-CLIP [22] 60.5 66.8 69.3 71.7 89.0 91.4 94.7 96.3 6.6 7.8 9.9 13.7 56.5
ViFi-CLIP [26] 63.0 65.1 69.6 72.0 91.0 93.7 95.0 96.4 6.7 7.9 10.2 13.5 57.0
TC-CLIP (Ours) 65.3 68.5 71.4 73.0 94.1 95.6 96.6 97.3 8.7 10.1 12.1 15.2 59.0
Table 9: Comparison with state-of-the-arts on base-to-novel generalization using Kinetics-400 pretrained model. All the models are first pretrained on Kinetics-400 and subsequently fine-tuned on each dataset.
HMDB-51 UCF-101 SSv2 All (Avg.)
Method Base Novel HM Base Novel HM Base Novel HM Base Novel HM
ActionCLIP [30] 69.0 57.2 62.6 85.6 75.3 80.1 8.1 8.7 8.4 54.2 47.1 50.4
A5 [13] 70.4 51.7 59.6 95.8 71.0 81.6 12.9 5.7 7.9 59.7 42.8 49.9
X-CLIP [22] 75.8 52.0 61.7 95.4 74.0 83.4 14.2 11.0 12.4 61.8 45.7 52.5
ViFi-CLIP [26] 77.1 54.9 64.1 95.9 74.1 83.6 15.8 11.5 13.3 62.9 46.8 53.7
TC-CLIP (Ours) 79.4 58.3 67.2 97.5 84.5 90.5 19.6 15.6 17.4 65.5 52.8 58.5
We additionally compare TC-CLIP with the state-of-the-art methods on few-shot and base-to-novel settings using the Kinetics-400 pretrained model. In this setting, all the models are first pretrained on the Kinetics-400 dataset and subsequently fine-tuned on each dataset. Table 8 summarizes the comparison results in the few-shot action recognition regime. Notably, TC-CLIP surpasses other methods by a large margin of at least 2%p in the average. Furthermore, Table 9 shows the comparison results in the base-to-novel generalization setup. In an average over all datasets, TC-CLIP shows distinct gaps from the baseline model ViFi-CLIP [26] by 3.5%p, 6%p, and 4.8%p in base, novel, and harmonic mean (HM), respectively.
Table 10: Temporal subset analysis using the temporal subset [27] on Kinetics-400 and SSv2. Gains over ViFi-CLIP are indicated in green.
K-400 fully-supervised SSv2 16-shot
Method All Temporal All Temporal
ViFi-CLIP [26] 83.9 87.8 12.4 25.9
TC-CLIP (Ours) 85.2 ($+$1.3) 89.2 ($+$1.4) 14.0 ($+$1.6) 29.9 ($+$4.0)
Table 11: VP layer selection ablation.
We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. The late-stage prompting design generalizes well.
$L_{\text{text}}$ $L_{\text{vision}}$ HMDB UCF SSv2 All
1 1 62.3 89.7 9.1 53.7
1 12 62.0 89.9 9.3 53.7
6 6 62.8 89.5 9.7 54.0
6 12 62.5 90.2 9.7 54.1
12 12 63.9 90.7 9.8 54.8
Table 12: VP initialization ablation by varying initialization cases of the VP layer and prompt vectors. We report $K$-shot training results where the top-1 accuracy is averaged over $K=2,4,8,16$. In general, it is advantageous to avoid random initialization.
Layer init. Prompt init. HMDB UCF SSv2 All
CLIP a photo of a 63.9 90.7 9.8 54.8
Random a photo of a 63.5 90.5 9.9 54.6
CLIP Random 62.5 90.2 9.5 54.1
Appendix 0.B Temporal Subset Analysis
We adopt the temporal subset analysis suggested in [27] to further analyze the temporal modeling ability of trained models. In this setting, the temporal subset consists of specific action classes that require more temporal information to recognize them. After shuffling frames in time and measuring whether the human annotator can still recognize the action, the classes that cannot be recognized are added to the temporal subset. The resulting temporal classes in Kinetics-400 and SSv2 are a total of 32 classes and 16 classes, respectively. Table 10 compares the top-1 accuracy using full validation splits and the temporal subsets. We use fully-supervised models for Kinetics-400 and 16-shot-trained models for SSv2. Our model shows unique gains over ViFi-CLIP [26] for the temporally challenging videos, more than when evaluated across the entire validation set. This demonstrates the superiority of TC-CLIP in handling temporal information.
Appendix 0.C More Ablation Study on VP
VP layer selection. Table 0.A presents the ablation study on the layer selection in the VP module. We compare the few-shot classification accuracy by varying the layer indices of the text and vision inputs $\{L_{\text{text}},L_{\text{vision}}\}$ in the VP module $f_{\theta_{\text{VP}}}(\mathbf{p}^{L_{\text{text}}-1},\mathbf{s}^{L_{\text{vision}}}_{\text{proj}})$. We observe that applying conditional prompting at the early stage ($L_{\text{text}}=1$) does not generalize well, regardless of the vision layer index. This suggests the early-stage prompting design is hard to generalize in a full fine-tuning scenario, possibly because CLIP was initially trained in a vision-text late-alignment fashion. The middle- and late-stage prompting designs are also compared, and the last layer indices are finally chosen for both modalities.
VP initialization. Table 0.A presents the ablation study on the initialization strategy of the VP layer and learnable prompt vectors. By default, we initialize the weight of the VP module using the weight from the last layer of the CLIP text encoder since randomly initializing it occasionally incurs unstable training results in the few-shot scenario. Similarly, it is beneficial to initialize the learnable prompt vectors using the prompt template “a photo of a” as generally done in the prompt tuning methods [37, 17].
Appendix 0.D Impact of Weight-space Ensemble
Figure 8: Weight averaging ablation.
In Fig. 8, we evaluate the effectiveness of weight ensembling by varying the ensemble ratio $w$ from 0 to 1 with a step size of 0.1. Specifically, the backbone weights of both vision and text encoders are linearly interpolated between CLIP and fine-tuned model, i.e., $\theta_{w}=(1-w)\cdot\theta_{\text{CLIP}}+w\cdot\theta_{\text{fine-tuned}}$. The $y$-axis displays the average accuracy on the zero-shot video datasets and the $x$-axis displays the accuracy on the fine-tuning dataset K-400. Our TC-CLIP achieves a better trade-off than the baseline as our curve is always on top of the baseline’s curve. This demonstrates that our model takes more advantages from weight ensembling. We choose $w=0.7$ as our final ensemble ratio.
Figure 9: Context token visualization of TC-CLIP on Kinetics-400, Kinetics-600, and SSv2 datasets. We visualize selected seed tokens and the resulting context tokens in the last layer of the vision encoder. Patch tokens with the same inner and border color are summarized into one context token. Regions highlighted in red represent a specific object or part grouped into a single context token throughout the video.
Figure 10: Attention visualization of TC-CLIP in comparison with ViFi-CLIP [26] on Kinetics-400, Kinetics-600, and SSv2 datasets. (a)–(b): TC-CLIP tends to focus more on fast-moving parts such as hands and arms. (c)–(d): While ViFi-CLIP dominantly attends to the most salient regions, TC-CLIP attends to multiple objects based on inter-object relationships relevant to the occurring actions. (e)–(f): TC-CLIP consistently attends to the main object with deformations throughout the video.
Appendix 0.E More Visualizations of Context Tokens and Attentions
Context token visualization. Fig. 9 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP. As can be seen, the seed tokens mainly consist of patch tokens from the most informative regions in each frame, which contain the foreground, such as a person, animals, hands, and objects. To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region. It is noteworthy that a single context token (highlighted in red) successfully tracks and summarizes a specific object or part throughout the entire video.
Attention visualization. Fig. 10 visualizes the attention maps of TC-CLIP compared to ViFi-CLIP [26]. As shown in Fig. 10(a)–(b), TC-CLIP tends to focus more on dynamically moving parts, such as hands and arms, during the action of throwing or shooting objects. Furthermore, TC-CLIP attends to multiple objects simultaneously based on inter-object relationships, as illustrated in Fig. 10(c)–(d). For example, during actions like “swinging baseball bat,” it effectively attends to both the bat and the baseball being struck. This highlights the TC-CLIP’s capability to understand complex interactions based on a temporal context in dynamic scenes, which is potentially more helpful than ViFi-CLIP, which only attends to the most salient region in each frame. Finally, Fig. 10(e)–(f) shows TC-CLIP’s consistent attention towards objects with deformations across video frames, which are more striking than ViFi-CLIP’s.
Appendix 0.F Experimental Setup Details
0.F.1 Dataset Details
We conduct experiments over 5 action recognition benchmarks: Kinetics-400 [16] & 600 [2], HMDB-51 [18], UCF-101 [28], and Something-Something v2 (SSv2) [8].
Kinetics-400 [16] is a large-scale action recognition dataset with a total of 400 action classes, where its video clips are collected from YouTube and last for about 10 seconds. It contains around 240k training videos and 20k validation videos.
Kinetics-600 [2] is an extension of Kinetics-400 with approximately 480k video clips covering 600 action categories. The videos are divided into 390k for training, 30k for validation, and 60k for testing. We mainly adopt the validation split for zero-shot evaluation.
HMDB-51 [18] dataset includes 6,869 clips divided into 51 action categories. There are three individual splits for training and validation.
UCF-101 [28] is an action recognition dataset collected from YouTube, including 13,320 video clips with 101 action categories. Similar to HMDB-51, the training and test videos have three splits.
SSv2 [8] is a challenging dataset with 174 fine-grained action classes, which are more temporally biased than the other datasets. The standard split consists of 168,913 training videos and 24,777 validation videos.
0.F.2 Implementation Details
During the bipartite matching, we start with the seed tokens arranged based on the [CLS] token attention values in each frame. These tokens are then divided into two sets by alternating positions. Subsequently, $r$ pairs of tokens with the highest cosine similarity are merged by averaging their features, and remaining two sets are then concatenated back together. This process is repeated iteratively, employing a constant $r$ scheduling for every iteration with an exception in the final iteration to ensure that the number of final context tokens becomes $k$. Following [37, 17], the learnable prompts are initialized with the prompt “a photo of a”. For training recipes, we follow [26] for zero-shot, few-shot, and fully-supervised settings and follow [9] for base-to-novel generalization. By default, we use the AdamW optimizer with momentum betas of (0.9, 0.98) and a weight decay of 0.001. The initial learning rate of the VP module is 10×10$\times$10 × larger than the base learning rate in each setting. Training configurations and evaluation metrics in each protocol are specified below.
Zero-shot action recognition. The models are trained on Kinetics-400 and evaluated on HMDB-51, UCF-101, and Kinetics-600 datasets. For HMDB-51 and UCF-101, we report the average and standard deviation of top-1 accuracy across three official validation splits. In the case of Kinetics-600, we apply the zero-shot evaluation protocol from [3], which exploits 220 categories of Kinetics-600 that do not appear in Kinetics-400. We use the three splits provided by [3], each containing 160 categories. The results include the average top-1 and top-5 accuracy and their respective standard deviations. During the training, the base learning rate is set to $8\times 10^{-6}$ and is decayed to $8\times 10^{-8}$ following the cosine decay scheduler. The batch size is 256, and the total number of epochs is 10, including 5 linear warmup epochs.
Few-shot action recognition. We adopt the $K$-shot training splits from [26] that randomly samples $K=2,4,8,16$ videos from each class on HMDB-51, UCF-101, and SSv2. The models are evaluated using the first validation split of HMDB-51 and UCF-101 and the full validation split of SSv2. The base learning rate is set to $2\times 10^{-6}$ and is decayed to $2\times 10^{-8}$. The batch size is 64, and the total number of epochs is set to 50, starting with 5 linear warmup epochs.
Base-to-novel generalization. We adopt the base and novel splits from [26]. The models are trained on a set of base (seen) classes in a few-shot manner and subsequently evaluated on a set of novel (unseen) classes for four datasets: Kinetics-400, HMDB-51, UCF-101, and SSv2. Each dataset comprises three training splits containing randomly sampled 16 shots of base action categories. We report the average accuracy over three splits. For HMDB-51 and UCF-101, the training and validation consider only their first split, whereas, for Kinetics and SSv2, the models are evaluated on their full validation split. The base learning rate is set to $3.33\times 10^{-6}$ and is decayed to $3.33\times 10^{-8}$. The batch size is 64. The number of epochs is 12, including 2 warmup epochs.
Fully-supervised action recognition. The models are trained on Kinetics-400 and evaluated on its complete validation split. The base learning rate is set to $2.2\times 10^{-5}$ and is decayed to $2.2\times 10^{-7}$ following the cosine decay scheduler. The batch size is 512, and the total epochs is 30 epochs, including 5 linear warmup epochs.

### v2

@@
 Abstract
-Pretrained vision-language models have shown effectiveness in video understanding.
-However, recent studies have not sufficiently leveraged essential temporal information from videos, simply averaging frame-wise representations or referencing consecutive frames.
-We introduce Temporally Contextualized CLIP (TC-CLIP), a pioneering framework for video understanding that effectively and efficiently leverages comprehensive video information.
-We propose Temporal Contextualization (TC), a novel layer-wise temporal information infusion mechanism for video that extracts core information from each frame, interconnects relevant information across the video to summarize into context tokens, and ultimately leverages the context tokens during the feature encoding process.
-Furthermore, our Video-conditional Prompting (VP) module manufactures context tokens to generate informative prompts in text modality.
-We conduct extensive experiments in zero-shot, few-shot, base-to-novel, and fully-supervised action recognition to validate the superiority of our TC-CLIP.
-Ablation studies for TC and VP guarantee our design choices.
-Code is available at https://github.com/naver-ai/tc-clip.
+We propose a novel framework for video understanding, called Temporally Contextualized CLIP (TC-CLIP), which leverages essential temporal information through global interactions in a spatio-temporal domain within a video.
+To be specific, we introduce Temporal Contextualization (TC), a layer-wise temporal information infusion mechanism for videos, which 1) extracts core information from each frame, 2) connects relevant information across frames for the summarization into context tokens, and 3) leverages the context tokens for feature encoding.
+Furthermore, the Video-conditional Prompting (VP) module processes context tokens to generate informative prompts in the text modality.
+Extensive experiments in zero-shot, few-shot, base-to-novel, and fully-supervised action recognition validate the effectiveness of our model.
+Ablation studies for TC and VP support our design choices.
+Our project page with the source code is available at https://github.com/naver-ai/tc-clip.
+Keywords:
+Video Action Recognition Vision-Language Model
 footnotetext: Work done during an internship at NAVER AI Lab.
@@
 1 Introduction
+Figure 1: Comparison of attention maps between various temporal modeling approaches.
+Both (a) and (b) fail to recognize actions in the latter frames, whereas (c) exhibits weak discriminability due to sparse attention on the background.
+In contrast, (d) our method successfully focuses on informative regions across all frames, leading to the accurate action recognition result.
 Figure 2: Temporal information learning methods.
-Prior works considered temporal cues during the encoding process via (a) cross-frame attention [22, 32] with [CLS] token interactions or (b) temporal window expansion [33] by adding adjacent frame tokens to key-value pairs.
-However, the former lacks patch-level interactions, while the latter limits the range of temporal locality.
-(c) joint space-time attention allows full interactions across all tokens, but it is costly and incurs suboptimal performance (see Fig.
-3).
-(d) Unlike prior approaches, our method aggregates pivotal tokens from a broader range yet efficiently for enhanced temporal integration into key-value pairs.
+Prior works consider temporal cues during the encoding process via (a) cross-frame attention [30, 43] with [CLS] token interactions or (b) temporal window expansion [44] by adding adjacent frame tokens to key-value pairs.
+However, the former lacks patch-level interactions, while the latter limits the range of temporal interactions.
+(c) Joint space-time attention allows full interactions across all tokens, but it is costly and suboptimal in practice (see Fig.
+3.) (d) Unlike prior approaches, our method aggregates pivotal tokens from a broader range yet efficiently for enhanced temporal integration into key-value pairs.
+Figure 3: Pitfall of joint space-time attention.
+(a) Extending CLIP’s temporal sequence length degrades attention quality, presumably because it was not trained on such long sequences.
+(b) We compare the action recognition performance in the few-shot setting on diverse datasets.
+All existing methods fall behind our method.
 Pretrained large-scale Vision-Language Models (VLMs) have shown remarkable generalization capability in video understanding and have emerged as promising tools even for zero-shot or open-vocabulary recognition tasks [33, 12, 50].
-However, pretraining task-specific models using video-text pairs pose significant challenges, primarily due to substantial computational costs as well as excessive expense for annotated video-text data [31, 34].
-Consequently, recent studies in video understanding [30, 22, 13, 32, 26, 33, 5, 9] have shifted their focus towards employing image-based VLMs such as CLIP [24] via fine-tuning—aligning video representations with text embeddings derived from category names.
-Despite the successful generalization of CLIP in video recognition, existing approaches fail to fully exploit temporal information in the video feature learning process.
-Specifically, the token interactions remain restrictive since they either rely on simple averaging of frame-wise representations [26], or attempt to gather information in the temporal axis only through class tokens [32, 22] as illustrated in Fig.
-1(a).
-Although VCLIP [33] incorporates patch-level details by bringing keys and values from neighborhood frames in its self-attention operation, as in Fig.
-1(b), its temporal scope is still limited to a local window.
-Such naïve computation for video representations makes models biased towards static information in their representation learning (e.g., objects and backgrounds) and hampers learning temporal information (e.g., motion and temporal variations).
+However, pretraining task-specific models using video-text pairs pose significant challenges, primarily due to substantial computational costs and excessive expense for annotated video-text data [42, 45].
+Consequently, recent studies in video understanding [41, 30, 15, 43, 35, 44, 5, 11] have shifted their focus toward employing image-based VLMs such as CLIP [33] with fine-tuning for aligning video representations with text embeddings derived from category names.
+Despite the success of CLIP in video recognition, existing approaches fail to model temporal information in the video feature learning process, as shown in Fig.
+1(a)-(b).
+This limitation stems from the restrictive token interactions in the temporal axis.
+For instance, cross-frame attention approaches [30, 43], shown in Fig.
+2(a), attempt to gather temporal information only through class tokens.
+Although VCLIP [44] incorporates patch-level details by bringing keys and values from neighborhood frames in its self-attention operation as in Fig.
+2(b), its temporal scope is too narrow.
+Furthermore, ViFi-CLIP [35] simply averages frame-wise representations with no inter-frame information exchanges.
+Such naïve approaches tend to bias the models towards static information in their representation learning (e.g., objects and backgrounds) and hamper learning temporal dynamics (e.g., motion and temporal variations).
 To ensure the global interactions of patch tokens in a spatio-temporal domain, one possible option is to consider every patch token from all frames as a reference during the encoding process as illustrated in Fig.
-1(c).
-Unfortunately, such a straightforward extension for temporally global interactions in VLMs pretrained with short image-text pairs witnesses extrapolation challenges [23, 4] due to the huge number of tokens in each video.
-We have observed that a naïve extension of sequence length along the temporal axis degrades its discriminability substantially, as shown in Fig.
-2.
-The joint space-time attention model distributes its attention broadly across the patches but fails to concentrate on informative tokens to recognize actions, incurring suboptimal performance compared to the frame-wise attention baseline.
-Moreover, this approach suffers from heavy computational overhead due to numerous redundant and similar tokens often corresponding to background regions.
-This paper presents Temporally Contextualized CLIP (TC-CLIP), a novel paradigm for extending CLIP to the video domain by encoding holistic video information through advanced temporal analysis.
-Specifically, our Temporal Contextualization (TC) pipeline summarizes global-range action cues into a small set of tokens for reference during the encoding process.
-Based on the simple idea by utilizing [CLS] query for self-attention, we could extract informative tokens from each frame and group them contextually.
-These context tokens act as additional key-value pairs for attention operations in the subsequent layer, presumably serving as a temporal bridge that conveys the video-level context throughout the encoding process.
+2(c).
+Unfortunately, such a straightforward extension for temporally global interactions in VLMs pretrained with short image-text pairs witnesses extrapolation challenges [32, 4]; we have observed that a naïve extension of sequence length along the temporal axis degrades its discriminability substantially, as shown in Fig.
+3(a).
+The joint space-time attention model spreads attention over many patches and fails to focus on informative tokens to recognize actions, resulting in suboptimal performance compared to the frame-wise attention baseline.
+Moreover, this approach suffers from heavy computational overhead due to numerous redundant and similar tokens, which often correspond to background regions.
+This paper presents Temporally Contextualized CLIP (TC-CLIP), a novel paradigm for extending CLIP to videos by encoding holistic video information through advanced temporal analysis.
+Specifically, our Temporal Contextualization (TC) pipeline summarizes global action cues into a small set of tokens, called context tokens, for reference during the encoding process.
+These context tokens act as additional key-value pairs for attention operations, presumably serving as temporal bridges that convey the video-level context.
 Our preliminary study, shown in Fig.
-3, implies that existing methods shown in Fig.
-1 offer minimal improvement over the frame-wise attention, highlighting the need for enhanced token interactions.
-Figure 2: Pitfall of joint space-time attention. Extending CLIP’s temporal sequence length degrades attention quality, presumably because it was not trained on such long sequences.
-Figure 3: Quantitative analysis on temporal modelings. We compare the action recognition performance in the few-shot settings on diverse datasets. All existing methods fall behind our method.
-Furthermore, we propose the Video-conditional Prompting (VP) module for the generation of instance-level textual prompts based on produced context tokens.
-The VP module is composed of cross-attention operations that adopt learnable text prompts as queries and visual temporal tokens as keys and values for injecting video instance representations into video-conditional textual prompts.
-This strategy alleviates the lack of textual semantics in video action recognition datasets, where category names are the only descriptions of actions (e.g., skateboarding, skydiving, ski jumping), without detailed narratives about acting subjects, observed objects and environment, etc.
-[…]
-2 Related Work
-CLIP for video action recognition.
-Recent video understanding studies [30, 13, 22, 32, 26, 33, 20, 5, 9] focus on adapting language-image pre-trained models [24, 10, 36].
-As a pioneering work, ActionCLIP [30] introduces CLIP [24] into video recognition by adding temporal attention layers after the backbone in a late-fusion fashion.
-To consider global-range temporal cues during feature encoding, X-CLIP [22] and Vita-CLIP [32] perform cross-frame attention, promoting [CLS] token interaction before spatial attention at each layer.
-On the other hand, there are lines of work that fully fine-tune CLIP with minimal architectural modifications [26, 33, 5, 9].
-ViFi-CLIP [26], a simple fine-tuning method with CLIP without additional temporal modules, generalizes well in open-vocabulary learning.
-Open-VCLIP [33] expands the temporal view of self-attention by adding keys and values of neighboring frames.
-Despite the generalization capacity of recent models, we argue they did not adequately leverage temporal information during the video feature encoding process, as they either rely on averaging frame-wise representations [26], [CLS] token interactions [22, 32], or considering only consecutive frames [33].
-Unlike prior approaches, we explore a temporal information infusion method to ensure spatio-temporal representation modeling in a broader range.
-Token aggregation.
-Conventional token aggregation methods, prioritizing efficiency enhancement, categorize into hard and soft clusterings.
-In hard clustering methods like K-Means [21], K-Medoids [15], and DPC-KNN [12], clusters are distinct without overlap, and every data point is assigned solely to one cluster.
-Bipartite matching [14] divides the data points into two groups and merges a specific number of edges with the highest scores.
-Soft clustering allows each data point to belong to multiple clusters with varying likelihood.
-Fuzzy C-Means [7] extends K-Means by enabling non-exclusive partitions and prototypes.
-Recent token aggregation methods for efficient ViTs [19, 1, 25, 35, 29] leverage aggregation algorithms to reduce the number of tokens for efficiency.
-While previous approaches focus on efficiency, we employ token aggregation approaches to link relevant tokens across the videos for contextual information.
-Prompt learning.
-Many works have proposed to transfer VLMs to downstream tasks by optimizing a discrete set of prompt vectors [11, 38, 37, 17, 26, 32, 13].
-In video recognition, [13] has introduced text prompt tuning, while ViFi-CLIP [26] and Vita-CLIP [32] perform prompting in both vision and text branches.
-However, these prompt vectors are individually optimized and not shared across the modalities.
-In image recognition, Co-CoOp [37] performs an instance-conditional prompt tuning by explicitly conditioning the text prompts on [CLS] tokens from image instances.
-MaPLe [17] learns multi-modal prompting by sharing layerwise context prompts for both branches.
-Unlikely, we generate video-conditional prompts by utilizing contextualized tokens as vision inputs and injecting summarized video information into text prompt vectors.
+3(b), implies that existing methods illustrated in Fig.
+2 offer minimal gains over the frame-wise attention, highlighting the need for enhanced token interactions.
+In addition, a Video-conditional Prompting (VP) module generates instance-level textual prompts based on context tokens from the vision encoder.
+The VP module comprises cross-attention operations that adopt learnable text prompts as queries and context tokens as keys and values to inject video instance representations into video-conditional textual prompts.
+This strategy compensates for the lack of textual semantics in action recognition datasets, where textual descriptions are limited to action class names (e.g., skateboarding, skydiving) without detailed narratives.
+To verify the effectiveness and robustness of TC-CLIP, we perform extensive evaluations across diverse video recognition benchmarks.
+Quantitative comparisons in zero-shot, few-shot, base-to-novel, and fully-supervised settings show that the proposed approach outperforms the state-of-the-art methods by significant margins.
+We also provide an in-depth analysis of our design choices and the impact of each component in our model.
 2 Proposed Method
@@
 In particular, we discuss the encoding procedure based on the vision and text encoders of CLIP, denoted by $\{f_{\theta_{v}},f_{\theta_{c}}\}$, to obtain video and text features, $\{\mathbf{v},\mathbf{c}\}$.
-3.1.1 Video encoding.
-Given a video $V\in\mathbb{R}^{T\times H\times W\times 3}$ of spatial resolution $H\times W$ with $T$ sampled frames, following the Vision Transformer (ViT) architecture [6], we first divide each frame into $P\times P$ non-overlapping patches and flattening them as a set of vectors $\{\mathbf{x}_{t,i}\in\mathbb{R}^{3P^{2}}\}_{i=1}^{N}$, where $t$ is the frame index, $i$ is the patch index, and $N=HW/P^{2}$ is the number of patches.
-A sequence of tokens in the $t^{\text{th}}$ frame, denoted by $\mathbf{z}_{t}^{0}$, is fed into the CLIP vision encoder after combining with the spatial positional encoding $\mathbf{e}_{\text{pos}}$ as
+Video encoding.
+Suppose that there exists a video $V\in\mathbb{R}^{T\times H\times W\times 3}$ of a spatial resolution $H\times W$ with $T$ sampled frames.
+Following the Vision Transformer (ViT) architecture [8], we first divide each frame into $P\times P$ non-overlapping patches and flatten them as a set of vectors $\{\mathbf{x}_{t,i}\in\mathbb{R}^{3P^{2}}\}_{i=1}^{N}$, where $t$ is the frame index, $i$ is the patch index, and $N=HW/P^{2}$ is the number of patches.
+Then, we derive a frame-level token sequence, $\mathbf{z}_{t}^{0}$ as follows:
 $\mathbf{z}_{t}^{0}=[\mathbf{x}_{\text{cls}},\mathbf{x}_{t,1}\mathbf{W}_{\text{emb}},\mathbf{x}_{t,2}\mathbf{W}_{\text{emb}},\ldots,\mathbf{x}_{t,N}\mathbf{W}_{\text{emb}}]+\mathbf{e}_{\text{pos}},$ (1)
-where $\mathbf{x}_{\text{cls}}\in\mathbb{R}^{d_{v}}$ is a learnable classification embedding, [CLS] token, and $\mathbf{W}_{\text{emb}}\in\mathbb{R}^{3P^{2}\times d_{v}}$ is a linear projection matrix.
-The CLIP vision encoder, $f_{\theta_{v}}(\cdot)$, sequentially encodes the frame-level representation $\mathbf{z}_{t}^{l}$ at each layer $l\in\{1,\ldots,L_{v}\}$ as follows:
-$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1}),$ (2)
-where $f_{\theta_{v}}^{l}(\cdot)$ is the $l^{\text{th}}$ layer of the CLIP vision encoder.
-Then, we project the [CLS] token of the $t^{\text{th}}$ frame, $\mathbf{z}_{t,0}^{L_{v}}$, onto a common vision-language latent space:
-$\mathbf{v}_{t}=\mathbf{z}_{t,0}^{L_{v}}\mathbf{W}_{\text{vis}},$ (3)
-where $\mathbf{W}_{\text{vis}}\in\mathbb{R}^{d_{v}\times d_{vl}}$ is a projection matrix.
-Finally, the video representation $\mathbf{v}$ is obtained by averaging the per-frame representations $\mathbf{v}_{t}$ as
-$\mathbf{v}=\text{AvgPool}([\mathbf{v}_{1},...,\mathbf{v}_{T}]).$ (4)
-3.1.2 Text encoding.
-Given a text description $C$ for a video, the input of the text encoder, $\textbf{c}^{0}$, is obtained by tokenizing words in the description and computing their word embedding vectors.
+where $\mathbf{x}_{\text{cls}}\in\mathbb{R}^{d_{v}}$ is a learnable classification embedding named [CLS] token, $\mathbf{W}_{\text{emb}}\in\mathbb{R}^{3P^{2}\times d_{v}}$ is a linear projection matrix, and $\mathbf{e}_{\text{pos}}$ is the spatial positional embedding.
+The CLIP vision encoder, $f_{\theta_{v}}(\cdot)$, sequentially encodes $\mathbf{z}_{t}^{l}$ at each layer $l\in\{1,\ldots,L_{v}\}$, which is given by
+$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1}).$ (2)
+We project the [CLS] token of the $t^{\text{th}}$ frame, denoted by $\mathbf{z}_{t,0}^{L_{v}}$, onto a common vision-language latent space using a matrix $\mathbf{W}_{\text{vis}}\in\mathbb{R}^{d_{v}\times d_{vl}}$, i.e., $\mathbf{v}_{t}=\mathbf{z}_{t,0}^{L_{v}}\mathbf{W}_{\text{vis}}$.
+Finally, the video representation $\mathbf{v}$ is obtained by averaging the per-frame representations $\mathbf{v}_{t}$ as $\mathbf{v}=\text{AvgPool}([\mathbf{v}_{1},...,\mathbf{v}_{T}])$.
+Text encoding.
+Given a text description $C$—category name in our problem—for a video, the input of the text encoder, $\textbf{c}^{0}$, is obtained by tokenizing words in the description and computing their word embedding vectors.
 In addition to the embeddings from category names, one can augment a sequence of prompt embeddings $\textbf{p}^{0}$, which are obtained from either hand-crafted templates such as “a photo of a” or learnable prompt vectors.
-The CLIP text encoder, $f_{\theta_{c}}(\cdot)$, sequentially processes a sequence of text embeddings including prompt embeddings, which is denoted by $[\textbf{p}^{0},\textbf{c}^{0}]$, and computes an intermediate representation at each layer $l\in\{1,\ldots,L_{c}\}$ as follows:
+The CLIP text encoder, $f_{\theta_{c}}(\cdot)$, sequentially processes a sequence of text embeddings including prompt embeddings, denoted by $[\textbf{p}^{0},\textbf{c}^{0}]$, and computes an intermediate representation at each layer $l\in\{1,\ldots,L_{c}\}$ as follows:
 $[\mathbf{p}^{l},\mathbf{c}^{l}]=f_{\theta_{c}}^{l}([\mathbf{p}^{l-1},\mathbf{c}^{l-1}]),$ (3)
@@
 The final text representations $\mathbf{c}$ is obtained by projecting the [EOS] token from the last layer to the vision-language latent space using a matrix $\mathbf{W}_{\text{text}}\in\mathbb{R}^{d_{l}\times d_{vl}}$, i.e., $\mathbf{c}=\mathbf{c}_{\text{eos}}^{L_{c}}\mathbf{W}_{\text{text}}$.
-3.1.3 Video-text alignment.
-Finally, the similarity between video and text embeddings are formulated as follows:
-$\mathrm{sim}(\mathbf{v},\mathbf{c})=\frac{\langle\mathbf{v},{\mathbf{c}\rangle}}{\lVert\mathbf{v}\rVert\lVert\mathbf{c}\rVert}.$ (6)
-During training, the goal is to maximize the similarity if $V$ and $C$ are matched but minimize it otherwise.
-For inference, the category with the highest similarity is chosen as the prediction output.
+Video-text alignment.
+The similarity between video and text embeddings are formulated as $\mathrm{sim}(\mathbf{v},\mathbf{c})=\frac{\langle\mathbf{v},{\mathbf{c}\rangle}}{\lVert\mathbf{v}\rVert\lVert\mathbf{c}\rVert}$.
+During training, the goal is to maximize the similarity if $V$ and $C$ are matched and minimize otherwise.
+For inference, the category with the highest similarity is chosen as the prediction.
+2.2 Motivation
+Table 1: Motivation: What is the proper format for reference tokens?
+We compare 16-shot training results using various types of reference tokens during the frame-level representation encoding process.
+Using context tokens consistently improves the baseline model regardless of the choice of the token aggregation function.
+Type of reference tokens HMDB-51 UCF-101 SSv2
+Baseline (No reference tokens) 67.1 93.3 12.0
+[CLS] tokens from all frames [30, 43] 67.2 (+0.1) 93.2 ($-0.1$) 12.3 (+0.3)
+Patch tokens from adjacent frames [44] 67.8 (+0.7) 93.2 ($-0.1$) 12.8 (+0.8)
+Patch tokens from all frames 63.3 ($-3.8$) 91.9 ($-1.4$) 12.0 (+0.0)
+Context tokens — K-means [26] 68.0 (+0.9) 93.3 (+0.0) 13.1 (+1.1)
+Context tokens — DPC-KNN [14] 67.9 (+0.8) 94.0 (+0.7) 14.3 (+2.3)
+Context tokens — Bipartite soft matching [16, 1] 68.0 (+0.9) 93.8 (+0.5) 14.3 (+2.3)
+Context tokens — Saliency-aware bipartite matching [6] 67.3 (+0.2) 93.7 (+0.4) 13.6 (+1.6)
+Despite the successful generalization of CLIP for action recognition, its visual feature encoding process in Eq.
+(2) constrains the model’s ability to capture comprehensive spatio-temporal dynamics because it only considers intra-frame token relationships.
+This limitation has led previous works to additionally incorporate reference tokens, denoted by $\mathbf{s}$, to encode the $t^{\text{th}}$ frame tokens $\mathbf{z}_{t}$ as
+$\mathbf{z}_{t}^{l}=f_{\theta_{v}}^{l}(\mathbf{z}_{t}^{l-1},\mathbf{s}^{l-1}).$ (4)
+However, their reference token designs are still limited due to insufficient spatio-temporal interaction range.
+For instance, cross-frame attention (Fig.
+2(a)) [30, 43] utilizes learnable global embedding vectors, e.g., [CLS] tokens, from all frames to define the reference token as $\mathbf{s}=[\mathbf{z}_{1,0},...,\mathbf{z}_{T,0}]$, and temporal window expansion (Fig.
+2(b)) [44], on the other hand, integrates neighboring frame patch tokens for $\mathbf{s}=[\mathbf{z}_{t-1},\mathbf{z}_{t+1}]$.
+Note that the former lacks patch-level details whereas the latter captures temporal information only within a local range.
+Although incorporating all patch tokens from a whole video as a reference (Fig.
+2(c)), $\mathbf{s}=[\mathbf{z}_{1},...,\mathbf{z}_{T}]$, might be conceptually reasonable, it is not practical due to the excessive number of tokens.
+Furthermore, this approach conflicts with the properties of CLIP pretrained on short image-text pairs and significantly degrades attention quality.
+To this end, we compute a reference, $\mathbf{s}=\phi([\mathbf{z}_{1},...,\mathbf{z}_{T}])$, using a small set of context tokens that summarize a whole input video, where $\phi(\cdot)$ is a token aggregation function.
+This approach delivers the essential temporal information for the feature encoding while maintaining CLIP’s effective sequence length.
+Our preliminary study, presented in Table 1, shows that using the proposed context tokens as a reference consistently outperforms the frame-wise attention baseline, whereas all other types of reference tokens only yield marginal gains or even suffer from performance degradation.
 2.3 Temporal Contextualization (TC)
-To alleviate the chronic drawbacks of existing CLIP-based video representations and leverage the spatio-temporal context in an entire video during the encoding process, we propose Temporal Contextualization (TC).
-TC consists of three steps, which are 1) informative token selection in each frame, 2) context summarization across spatio-temporal dimensions, and 3) temporal context infusion to all tokens in the subsequent layer.
+Figure 4: Overview of Temporal Contextualization (TC).
+We first collect informative tokens from each frame and then aggregate relevant seed tokens to obtain context tokens.
+They are used as key-value pairs for the self-attention in the next layer.
+Based on our motivation, we propose Temporal Contextualization (TC), which consists of three steps: 1) informative token selection in each frame, 2) context summarization across spatio-temporal dimensions, and 3) context infusion to all tokens in the subsequent layer.
 Fig.
 4 illustrates an overview of TC.
-We discuss the three steps of our TC pipeline in the following.
-Figure 4: Overview of Temporal Contextualization (TC).
-(a) Starting with the original patch tokens of the entire video, we first select informative seed tokens from each frame and perform temporal context summarization to aggregate relevant tokens to obtain context tokens.
-(b) The encoded context tokens are subsequently utilized as key-value pairs for self-attention in the next layer.
-(c) We employ attention score with [CLS] tokens as the criterion for seed token selection.
-3.2.1 Informative token selection.
-Due to a large number of redundant patches, the use of all patches for encoding a video may be suboptimal for extracting desired temporal information.
-Our approach focuses on selecting the informative tokens before performing the token aggregation.
-To this end, the attention scores obtained from self-attention operations in each frame are used as criteria to select seed tokens.
-Specifically, given patch tokens $\{\mathbf{z}_{t,i}\}_{i=1}^{N}$ in the $t^{\text{th}}$ frame, a set of attention scores $\{\mathbf{a}(\mathbf{z}_{t,i})\}_{i=1}^{N}$ is driven from the attention of the query of the [CLS] token to the keys of patch tokens, which is given by
+Informative token selection.
+Due to the many redundant tokens in a video, using all patches may be suboptimal for extracting desired temporal information.
+We only select the informative seed tokens based on each frame’s attention scores obtained from self-attention operations.
+Specifically, given patch tokens $\{\mathbf{z}_{t,i}\}_{i=1}^{N}$ in the $t^{\text{th}}$ frame, a set of attention scores $\{\mathbf{a}(\mathbf{z}_{t,i})\}_{i=1}^{N}$ is driven from the attentiveness of the [CLS] token with respect to other tokens, which is given by
 $\mathbf{a}(\mathbf{z}_{t})=\text{Softmax}\Big{(}\frac{\mathbf{q}_{\text{cls}}\mathbf{K}_{\mathbf{z}_{t}}^{\mathsf{T}}}{\sqrt{d}}\Big{)},$ (5)
 where both the query $\mathbf{q}_{\text{cls}}=\mathbf{z}_{\text{t,0}}\mathbf{W}_{q}\in\mathbb{R}^{d}$ and keys $\mathbf{K}_{\mathbf{z}_{t}}=\mathbf{z}_{t}\mathbf{W}_{k}\in\mathbb{R}^{(N+1)\times d}$ are given by linear projections of input $\mathbf{z}_{t}\in\mathbb{R}^{(N+1)\times d}$.
-In practice, our model yields multiple [CLS] attention scores through multi-head self-attention (MHSA) operations and compute the average of the attention scores from all heads, i.e., $\bar{\mathbf{a}}_{t,i}=\sum_{h=1}^{H}\mathbf{a}_{t,i}^{h}/H$, where $\mathbf{a}_{t,i}^{h}=\mathbf{a}^{h}(\mathbf{z}_{t,i})$ is the attention score for the $i^{\text{th}}$ patch $\mathbf{z}_{t,i}$ in the $t^{\text{th}}$ frame and $H$ is a number of heads.
-Finally, we formulate a set of seed tokens for the $t^{\text{th}}$ frame, $\mathcal{S}_{t}$, by selecting the indices for $n_{s}$ elements with highest attention score $\bar{\mathbf{a}}_{t,i}$, where the number of seed tokens in each frame $n_{s}$ is controlled by the ratio of $\alpha=n_{s}/N$.
-3.2.2 Temporal context summarization.
-We now interconnect the seed tokens derived from individual frames based on their relevance and identify a collection of context tokens.
-Specifically, we collect all the seed tokens from all frames as $\mathcal{S}=\{(t,i)|i\in\mathcal{S}_{t},t=1,\ldots,T\}$ and perform their spatio-temporal summarization by clustering and merging all the seed tokens in $\mathcal{S}$ as follows:
-$\hat{\mathbf{s}}=\phi\big{(}\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}\big{)}.$ (8)
-where $\hat{\mathbf{s}}\in\mathbb{R}^{k\times D}$ denotes a collection of the summarized tokens, which we call context tokens, $\phi$ is a token aggregation function, and $\hat{\textbf{z}}_{t,i}$ indicates an interim token encoded from $\textbf{z}_{t,i}$ via the self-attention operation.
-In practical, we construct a bipartite graph whose vertices and edges represent the original seed tokens and their pairwise similarities.
-Then, we iteratively perform bipartite matching [14] that splits the vertex set and merges $r$ pairs of vertices based on their similarities by simple averaging of vertex features until there remain $k$ number of tokens.
-Subsequently, the context tokens $\hat{\mathbf{s}}$ are fed into a feed-forward network (FFN) as done in the original encoding procedure and are forwarded to the next layer.
-3.2.3 Temporal context infusion.
+In practice, our model yields multiple [CLS] attention scores from multi-head self-attention (MHSA) operations and computes the average of the attention scores from all heads, i.e., $\bar{\mathbf{a}}_{t,i}=\sum_{h=1}^{H}\mathbf{a}_{t,i}^{h}/H$, where $\mathbf{a}_{t,i}^{h}=\mathbf{a}^{h}(\mathbf{z}_{t,i})$ is the attention score for the $i^{\text{th}}$ patch $\mathbf{z}_{t,i}$ in the $t^{\text{th}}$ frame and $H$ is the number of heads.
+Finally, we identify a set of seed token indices for the $t^{\text{th}}$ frame, $\mathcal{S}_{t}$, by selecting $n_{s}$ elements with the highest attention scores, where $n_{s}$ is controlled by a hyperparameter $\alpha=n_{s}/N$.
+Temporal context summarization.
+We describe how to connect the seed tokens derived from individual frames based on their relevance and identify a collection of context tokens.
+We first collect the seed tokens from all frames as $\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}$, where $\mathcal{S}=\{(t,i)|i\in\mathcal{S}_{t},t=1,\ldots,T\}$ is a set of seed token indices from all frames and $\hat{\textbf{z}}_{t,i}$ indicates an interim token encoded from $\textbf{z}_{t,i}$ via the self-attention operation.
+Then we perform their spatio-temporal summarization by clustering and merging all the seed tokens as
+$\hat{\mathbf{s}}=\phi\big{(}\{\hat{\mathbf{z}}_{t,i}\}_{(t,i)\in\mathcal{S}}\big{)},$ (6)
+where $\hat{\mathbf{s}}\in\mathbb{R}^{k\times D}$ denotes a collection of the summarized tokens, which we call context tokens, and $\phi$ is a token aggregation function.
+While diverse token aggregation techniques are valid for TC (See Table 8), we adopt bipartite soft matching [16, 1] by default.
+Subsequently, the context tokens $\hat{\mathbf{s}}$ are fed into a feed-forward network (FFN).
+Temporal context infusion.
 Finally, we infuse the summarized context to all patch tokens by modifying the self-attention function.
@@
 where $\mathbf{K}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{k}$ and $\mathbf{V}_{\mathbf{s}}=\mathbf{s}\mathbf{W}_{v}$ are linear projections of the context tokens $\mathbf{s}\in\mathbb{R}^{k\times d}$.
-Here, $\mathbf{B}\in\mathbb{R}^{(N+1)\times(N+k+1)}$ is a bias matrix that distinguishes between frame-level local information and video-level global information in the expanded key matrix as
+Here, $\mathbf{B}\in\mathbb{R}^{(N+1)\times(N+k+1)}$ is a bias matrix that distinguishes between frame-level local information and video-level global information in the expanded key matrix as follows:
 $\mathbf{B}_{ij}=\begin{cases}b_{\text{local}}&\text{if }j\leq N+1\\ b_{\text{global}}&\text{otherwise},\end{cases}$ (8)
@@
 $\mathbf{s}^{l}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{s}}^{l}))+\hat{\mathbf{s}}^{l},$ (11)
-where $\text{MHSA}_{\text{TC}}(\cdot,\cdot)$ denotes the TC-based MHSA operation including Eq. (9) and $\text{LN}(\cdot)$ stands for the layer normalization function.
+where $\text{MHSA}_{\text{TC}}(\cdot,\cdot)$ denotes the MHSA operation based on Eq. (7) and $\text{LN}(\cdot)$ stands for the layer normalization function.
 2.4 Video-conditional Prompting (VP)
-We present the Video-conditional Prompting (VP) module, which aims to further leverage the precisely aggregated temporal information in visual domain for text encoding.
-We impose a cross-attention between prompt vectors and context tokens to enrich the information in the prompt vectors, and the detailed procedure is illustrated in Fig.
+Figure 5: Video-conditional Prompting (VP) module.
+Video information from the context tokens is injected into the text prompt vectors using a cross-attention mechanism, generating instance-level prompts that make up for the lack of textual semantics.
+The Video-conditional Prompting (VP) module further leverages the comprehensive video information derived from the visual domain for text encoding.
+We apply a cross-attention between prompt vectors and context tokens to enrich the information in the prompt vectors as illustrated in Fig.
 5.
 Let $\textbf{c}^{l-1}$ and $\textbf{p}^{l-1}$ be class name tokens and learnable prompt vectors from the $(l-1)^{\text{th}}$ layer of the text encoder, respectively.
-We derive temporally contextualized prompt vectors $\hat{\mathbf{p}}^{l-1}$ by passing the layer-normalized prompt tokens and context tokens through a cross-attention layer as follows:
-Figure 5: Video-conditional Prompting (VP) module.
-Video information from the context tokens is injected to the text prompt vectors based on a cross-attention mechanism, generating instance-level prompts that support the lack of textual semantics.
+We derive temporally contextualized prompt vectors $\tilde{\mathbf{p}}^{l-1}$ by passing the layer-normalized prompt tokens and context tokens through a cross-attention layer and an FFN layer as follows:
 $\mathbf{s}_{\text{proj}}^{l}$ $=\text{SG}(\mathbf{s}^{l}\mathbf{W}_{\text{vis}}),$ (12)
-$\hat{\mathbf{p}}^{l-1}$ $=\text{MHCA}(\text{LN}_{p}(\mathbf{p}^{l-1}),\text{LN}_{s}(\mathbf{s}_{\text{proj}}^{l}))+\mathbf{p}^{l-1}$ (15)
-where $\text{SG}(\cdot)$ is a stop-gradient function, $\mathbf{W}_{\text{vis}}$ is a weight matrix of CLIP to linearly project vision representations to a common vision-language latent space, and $\text{MHCA}(\cdot,\cdot)$ is a multi-head cross-attention opearation for interactions across modalities, accepting text prompt vectors as queries and vision features as keys and values.
-Finally, the contextualized prompt vector $\hat{\mathbf{p}}^{l-1}$ are updated by FFN:
-$\tilde{\mathbf{p}}^{l-1}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{p}}^{l-1})+\hat{\mathbf{p}}^{l-1}$ (16)
+$\hat{\mathbf{p}}^{l-1}$ $=\text{MHCA}(\text{LN}_{p}(\mathbf{p}^{l-1}),\text{LN}_{s}(\mathbf{s}_{\text{proj}}^{l}))+\mathbf{p}^{l-1},$ (13)
+$\tilde{\mathbf{p}}^{l-1}$ $=\text{FFN}(\text{LN}(\hat{\mathbf{p}}^{l-1}))+\hat{\mathbf{p}}^{l-1},$ (14)
+where $\text{SG}(\cdot)$ is a stop-gradient function, $\mathbf{W}_{\text{vis}}$ is a weight matrix of CLIP to linearly project vision representations onto a common vision-language latent space, and $\text{MHCA}(\cdot,\cdot)$ is a multi-head cross-attention operation for interactions across modalities, accepting text prompt vectors as queries and vision features as keys and values.
 The VP module $f_{\theta_{\text{VP}}}(\cdot,\cdot)$ is defined by a composition of Eq.
@@
 $[\mathbf{p}^{l},\mathbf{c}^{l}]=\begin{cases}f_{\theta_{c}}^{l}([f_{\theta_{\text{VP}}}(\mathbf{p}^{l-1},\mathbf{s}_{\text{proj}}^{l}),\mathbf{c}^{l-1}])&\text{if}~{}l=L_{c}\\ f_{\theta_{c}}^{l}([\mathbf{p}^{l-1},\mathbf{c}^{l-1}])&\text{otherwise.}\end{cases}$ (15)
-The resulting instance-conditioned prompt vectors effectively support the lack of textual semantics in action recognition datasets.
-The impact of our VP design is verified in the experiments in Sec.
-5.
 2.5 Training Objective
-TC-CLIP learns to maximize the similarity of video representations $\mathbf{v}$ and text representations $\mathbf{c}$ for matching pairs in a mini-batch via the cross-entropy loss:
-$\mathcal{L}=-\sum_{i}\log\frac{\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{j})/\tau)},$ (18)
-where $\tau$ is a learnable temperature parameter. Our model is fully fine-tuned in an end-to-end manner.
+TC-CLIP learns to maximize the similarity of video representations $\mathbf{v}$ and text representations $\mathbf{c}$ for matching pairs in a mini-batch via the cross-entropy loss as $\mathcal{L}=-\sum_{i}\log\frac{\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{i})/\tau)}{\sum_{j}\text{exp}(\text{sim}(\mathbf{v}_{i},\mathbf{c}_{j})/\tau)},$ where $\tau$ is a learnable temperature parameter.
+Our model is fully fine-tuned in an end-to-end manner.
 3 Experiments
-4.1 Experimental Setup
-Datasets.
-We conduct experiments over 5 video benchmarks: Kinetics-400 [16] & 600 [2], HMDB-51 [18], UCF-101 [28], and Something-Something v2 (SSv2) [8].
-Following [26], our evaluation protocol includes zero-shot, few-shot, base-to-novel generalization, and fully-supervised action recognition.
-Implementation details.
-We adopt CLIP with ViT-B/16 for all experiments.
-Our baseline is ViFi-CLIP [26].
-During the training, we sample 16 frames to form a video clip.
-During the evaluation, two temporal clips with one spatial crop (2 $\times$ 1 view) per video are sampled to produce a prediction unless otherwise stated.
-For training recipes, we follow [26] for zero-shot, few-shot, and fully-supervised settings and follow [9] for base-to-novel generalization.
-Following [33, 20, 5, 9], we adopt linear weight-space ensembling of our fine-tuned model and CLIP for the zero-shot protocol.
-We set the number of merged tokens $r$ in each bipartite matching iteration to 100, and the weight of the VP module is initialized with the weight from the last layer of the CLIP text encoder.
-All models are trained using 4 NVIDIA Tesla V100 GPUs.
-See supplementary material for more details.
 Table 2: Comparison with state-of-the-arts on zero-shot action recognition.
 All the models are trained on Kinetics-400 and directly evaluated on other datasets.
-† denotes the methods with weight-space ensemble.
+WE indicates the weight-space ensemble between the fine-tuned model and CLIP, adopted for all applicable models for fair comparisons.
+† denotes results reproduced using our implementation.
+The best results are in bold-faced numbers, and the second-best ones are underlined.
 Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
-Method HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
+Method WE HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
 Vanilla CLIP [33] 40.8 $\pm$ 0.3 63.2 $\pm$ 0.2 59.8 $\pm$ 0.3 83.5 $\pm$ 0.2 54.6
-ActionCLIP [30] 40.8 $\pm$ 5.4 58.3 $\pm$ 3.4 66.7 $\pm$ 1.1 91.6 $\pm$ 0.3 55.3
+ActionCLIP [41] 49.1 $\pm$ 0.4 68.0 $\pm$ 0.9 56.1 $\pm$ 0.9 83.2 $\pm$ 0.2 57.7
 A5 [15] 44.3 $\pm$ 2.2 69.3 $\pm$ 4.2 55.8 $\pm$ 0.7 81.4 $\pm$ 0.3 56.5
@@
 Vita-CLIP [43] 48.6 $\pm$ 0.6 75.0 $\pm$ 0.6 67.4 $\pm$ 0.5 - 63.7
-ViFi-CLIP [26] 51.3 $\pm$ 0.6 76.8 $\pm$ 0.7 71.2 $\pm$ 1.0 92.2 $\pm$ 0.3 66.4
-Open-VCLIP [33] † 53.9 $\pm$ 1.2 83.4 $\pm$ 1.2 73.0 $\pm$ 0.8 93.2 $\pm$ 0.1 70.1
-TC-CLIP (Ours) † 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 94.4 $\pm$ 0.2 71.0
+ViFi-CLIP [35] 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 92.1 $\pm$ 0.3 67.3
+TC-CLIP (Ours) 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 93.2 $\pm$ 0.2 68.9
+ActionCLIP [41] 51.9 $\pm$ 0.5 74.2 $\pm$ 1.0 67.5 $\pm$ 1.2 90.7 $\pm$ 0.1 64.5
+ViFi-CLIP [35] 52.2 $\pm$ 0.7 81.0 $\pm$ 0.9 73.9 $\pm$ 0.5 93.3 $\pm$ 0.3 69.0
+Open-VCLIP [44] 53.9 $\pm$ 1.2 83.4 $\pm$ 1.2 73.0 $\pm$ 0.8 93.2 $\pm$ 0.1 70.1
+TC-CLIP (Ours) 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 94.4 $\pm$ 0.2 71.0
 Using LLM-based text augmentation
-MAXI [20] † 52.3 $\pm$ 0.7 78.2 $\pm$ 0.8 71.5 $\pm$ 0.8 92.5 $\pm$ 0.4 67.3
-OST [5] † 55.9 $\pm$ 1.2 79.7 $\pm$ 1.1 75.1 $\pm$ 0.6 94.6 $\pm$ 0.2 70.2
-FROSTER [9] † 54.8 $\pm$ 1.3 84.8 $\pm$ 1.1 74.8 $\pm$ 0.9 - 71.5
-TC-CLIP (Ours) † 56.0 $\pm$ 0.3 85.4 $\pm$ 0.8 78.1 $\pm$ 1.0 95.7 $\pm$ 0.3 73.2
-4.2 Quantitative Comparison
-We mainly compare our method with CLIP-based video recognition models: Vanilla CLIP [24], ActionCLIP [30], A5 [13], X-CLIP [22], Vita-CLIP [32], ViFi-CLIP [26], Open-VCLIP [33], OST [5], and FROSTER [9].
-To fairly compare our method with approaches that utilize Large Language Model (LLM)-based text augmentation [20, 5, 9], we divide our results into two versions: one using the original action category names (colored with blue) and the other adopting the LLM-rephrased category names from FROSTER [9] (colored with purple).
-Note that experiments on the SSv2 dataset do not involve LLM-rephrasing.
-[…]
+MAXI [25] 52.3 $\pm$ 0.7 78.2 $\pm$ 0.8 71.5 $\pm$ 0.8 92.5 $\pm$ 0.4 67.3
+OST [5] 55.9 $\pm$ 1.2 79.7 $\pm$ 1.1 75.1 $\pm$ 0.6 94.6 $\pm$ 0.2 70.2
+FROSTER [11] 54.8 $\pm$ 1.3 84.8 $\pm$ 1.1 74.8 $\pm$ 0.9 - 71.5
+TC-CLIP (Ours) 56.0 $\pm$ 0.3 85.4 $\pm$ 0.8 78.1 $\pm$ 1.0 95.7 $\pm$ 0.3 73.2
 Table 3: Comparison with state-of-the-arts on few-shot action recognition.
@@
 TC-CLIP (Ours) 58.6 63.3 65.5 68.8 86.8 90.1 92.0 94.3 7.3 8.6 9.3 14.0 54.9
-Table 3: Comparison with state-of-the-arts on base-to-novel generalization.
-All the models are directly fine-tuned from CLIP.
-† denotes the results are taken from [9].
-Our results using the original and LLM-rephrased category names are highlighted in blue and purple, respectively.
+Table 4: Comparison with state-of-the-arts on base-to-novel generalization. All the models are directly fine-tuned from CLIP. † results are taken from [11].
 K-400 HMDB-51 UCF-101 SSv2 All (Avg.)
@@
 ViFi-CLIP [35] 76.4 61.1 67.9 73.8 53.3 61.9 92.9 67.7 78.3 16.2 12.1 13.9 64.8 48.6 55.5
-Open-VCLIP [33] † 76.5 62.6 68.9 70.3 50.4 58.9 94.8 77.5 85.3 16.0 11.0 13.0 64.4 50.4 56.5
+Open-VCLIP [44] 76.5 62.6 68.9 70.3 50.4 58.9 94.8 77.5 85.3 16.0 11.0 13.0 64.4 50.4 56.5
 TC-CLIP (Ours) 78.9 63.6 70.4 73.3 54.1 62.2 95.5 78.0 85.9 17.5 13.4 15.2 66.3 52.3 58.5
@@
 TC-CLIP (Ours) 79.1 65.4 71.6 73.3 59.1 65.5 95.4 81.6 88.0 17.5 13.4 15.2 66.3 54.9 60.1
+Table 5: Fully-supervised action recognition results on Kinetics-400. Views means (temporal clips) $\times$ (spatial crops), and F denotes number of frames.
+Method Top-1 Top-5 F Views
+ActionCLIP [41] 83.8 96.2 32 10 $\times$ 3
+X-CLIP [30] 84.7 96.8 16 4 $\times$ 3
+Vita-CLIP [43] 82.9 96.3 16 4 $\times$ 3
+ViFi-CLIP [35] 83.9 96.3 16 4 $\times$ 3
+OST [5] 83.2 - 16 1 $\times$ 1
+TC-CLIP (Ours) 85.2 96.9 16 4 $\times$ 3
+[…]
+Method Zero Few B2N Full Params GFLOPs TP
+ActionCLIP [41] 64.5 46.5 46.0 83.8 143.7M 567 20
+X-CLIP [30] 60.6 50.2 49.5 84.7 169.7M 288 36
+Vita-CLIP [43] 63.7 - - 82.9 161.8M 307 30
+ViFi-CLIP [35] 69.0 52.9 55.5 83.9 124.3M 285 38
+Open-VCLIP [44] 70.1 - 56.5 - 124.3M 308 29
+TC-CLIP (Ours) 71.0 54.8 58.5 85.2 127.5M 304 24
+TC-CLIP (Ours)§ 70.7 54.4 58.6 84.9 127.5M 291 34
+We conduct experiments on 5 video benchmarks: Kinetics-400 [17] & 600 [2], HMDB-51 [22], UCF-101 [39], and Something-Something v2 (SSv2) [10].
+Following [35], our evaluation protocols include zero-shot, few-shot, base-to-novel generalization, and fully-supervised action recognition tasks.
+We adopt CLIP with ViT-B/16 for all experiments and our baseline is ViFi-CLIP [35].
+All models are trained using 4 NVIDIA Tesla V100 GPUs.
+More details are in the appendix.
+3.1 Quantitative Comparison
+We mainly compare our method with CLIP-based video recognition models: Vanilla CLIP [33], ActionCLIP [41], A5 [15], X-CLIP [30], Vita-CLIP [43], ViFi-CLIP [35], Open-VCLIP [44], OST [5], and FROSTER [11].
+For the fair comparisons with approaches based on Large Language Model (LLM) with text augmentation [25, 5, 11], we produce two versions of our results: one using the original action category names (colored in blue) and the other adopting the LLM-rephrased category names obtained from FROSTER [11] (colored in purple).
+Note that experiments on the SSv2 dataset do not involve LLM-rephrasing.
+Zero-shot action recognition.
+Table 2 exhibits the zero-shot generalization ability of several models, where they are trained on K-400 and then directly evaluated on individual datasets.
+For fair comparisons with recent models [44, 25, 5, 11], we employ weight-space ensembling (WE) for all applicable models except those freezing a backbone during fine-tuning.
+Specifically, the weights of both vision and text encoders are linearly interpolated between CLIP and the fine-tuned model as $\theta_{w}=(1-w)\cdot\theta_{\text{CLIP}}+w\cdot\theta_{\text{fine-tuned}}$.
+TC-CLIP consistently outperforms others across all datasets, showing its superior generalization ability.
 Few-shot action recognition.
-We demonstrate the learning capacity of our method under a challenging few-shot scenario.
-In this setting, models are directly fine-tuned from CLIP on HMDB-51, UCF-101, and SSv2 using $K$-shot samples, where $K$ is set to 2, 4, 8, and 16.
-Table 2 shows the comparison with state-of-the-art video action recognition methods on the few-shot setting.
-Our method achieves state-of-the-art performance with an average accuracy of 54.8%, surpassing our baseline model ViFi-CLIP [26] by 1.9%p.
+We verify the learning capacity of our method under a challenging few-shot scenario.
+In Table 3, models are directly fine-tuned from CLIP on each dataset using $K$-shot samples, where $K$ is 2, 4, 8, and 16.
+TC-CLIP achieves the best performance with large margins from ViFi-CLIP [35].
 Base-to-novel generalization.
-Similar to the few-shot setting, models are directly fine-tuned from CLIP on the base classes of K-400, HMDB-51, UCF-101, and SSv2, and evaluated for both base and novel classes.
-The top-1 accuracy on base and novel classes, and their harmonic mean (HM) are reported.
-Table 3 shows that our model achieves state-of-the-art performance on the novel classes and HM across all datasets.
-In particular, TC-CLIP achieves an average HM accuracy of 58.5%, surpassing previous state-of-the-arts by 2.0%p.
-Table 4: Comparison with state-of-the-arts on fully-supervised action recognition on Kinetics-400.
-Views are denoted with temporal clips $\times$ spatial crops.
-† means that computational costs are measured in our environment using a single V100 GPU.
-The results of other methods are taken from [26].
-Method Top-1 Top-5 Frames Views GFLOPs Params (M)
-ActionCLIP [30] 83.8 96.2 32 10 $\times$ 3 563 168.5
-X-CLIP [22] 84.7 96.8 16 4 $\times$ 3 287 131.5
-Vita-CLIP [22] † 82.9 96.3 16 4 $\times$ 3 307 185.3
-ViFi-CLIP [26] 83.9 96.3 16 4 $\times$ 3 281 124.7
-OST [5] 83.2 - 16 1 $\times$ 1 - -
-TC-CLIP (Ours) 85.2 96.9 16 4 $\times$ 3 304 127.5
+Similarly, models are directly fine-tuned from CLIP using the base classes of each dataset and evaluated for both base and novel classes.
+Table 4 reports top-1 accuracies on the base and novel classes with their harmonic mean (HM).
+TC-CLIP performs the best on the novel classes and HM across all datasets, especially showing solid results on the SSv2 dataset.
 Fully-supervised action recognition.
-Table 4 shows performance comparison results under fully-supervised setting, where the models are both trained and evaluated on the K-400 dataset.
-Our model achieves top-1 accuracy of 85.2% in the K-400 validation split, improving 1.3%p over our baseline ViFi-CLIP [26].
-Table 5: Component-wise ablation on the zero-shot setting. ($\Delta$) denotes the average top-1 accuracy improvement over baseline.
-Case TC VP WE HMDB-51 UCF-101 K-600 All ($\Delta$)
-Baseline 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 67.3
-(a) w/o Video-conditional Prompting (VP) 54.3 $\pm$ 0.6 81.9 $\pm$ 1.0 75.5 $\pm$ 0.5 70.6 ($+$3.3)
-(b) w/o Temporal Contextualization (TC) 53.2 $\pm$ 0.8 82.0 $\pm$ 0.9 74.7 $\pm$ 0.7 70.0 ($+$2.7)
-(c) w/o Weight-space Ensembling (WE) 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 68.9 ($+$1.6)
-TC-CLIP 54.2 $\pm$ 0.7 82.9 $\pm$ 0.6 75.8 $\pm$ 0.5 71.0 ($+$3.7)
-Table 6: Temporal Contextualization (TC) ablation on the few-shot setting.
-SE, JE, and Bias denote spatial embedding, joint space-time embedding, and bias (Eq.
-10).
+Table 5 shows performance comparison results under the fully-supervised setting, where the models are trained and evaluated both on the K-400 dataset.
+TC-CLIP achieves top-1 accuracy of 85.2% in the validation split, improving 1.3%p over our baseline ViFi-CLIP [35].
+[…]
+Table 7: Component-wise ablations on the zero-shot setting. $\Delta$ denotes the average top-1 accuracy gain over baseline.
+Without weight-space ensembling With weight-space ensembling
+Case HMDB-51 UCF-101 K-600 All ($\Delta$) HMDB-51 UCF-101 K-600 All ($\Delta$)
+Baseline 52.3 $\pm$ 0.2 78.9 $\pm$ 1.1 70.7 $\pm$ 0.8 67.3 52.2 $\pm$ 0.7 81.0 $\pm$ 0.9 73.9 $\pm$ 0.5 69.0
+(a) +TC 53.6 $\pm$ 0.2 78.6 $\pm$ 1.0 71.8 $\pm$ 0.7 68.0 (+0.7) 54.3 $\pm$ 0.6 81.9 $\pm$ 1.0 75.5 $\pm$ 1.0 70.6 (+1.6)
+(b) +VP 53.2 $\pm$ 0.8 80.5 $\pm$ 0.7 71.6 $\pm$ 0.9 68.4 (+1.1) 53.4 $\pm$ 0.8 82.0 $\pm$ 0.9 74.7 $\pm$ 0.7 70.0 (+1.0)
+(c) +TC+VP 53.7 $\pm$ 0.7 80.4 $\pm$ 0.9 72.7 $\pm$ 0.5 68.9 (+1.6) 54.2 $\pm$ 1.1 82.9 $\pm$ 0.9 75.8 $\pm$ 0.4 71.0 (+2.0)
+Table 8: Effect of TC with various token aggregation strategies.
+TC consistently outperforms the frame-wise attention baseline across several different token selection and merging methods.
+$K$-shot action recognition results are reported with the top-1 accuracy averaged over $K=2,4,8,16$.
 Default settings are marked in gray.
-(a) Seed token selection metric.
+(a) Seed token selection strategy.
+Case HMDB UCF SSv2 All ($\Delta$)
+Baseline 62.6 89.2 8.7 53.5
+No selection 62.8 89.8 9.7 54.1 (+0.6)
+Head-wise key norm 62.3 89.8 9.8 54.0 (+0.5)
+Averaged key norm 62.5 89.4 9.3 53.7 (+0.2)
+Head-wise CLS attn. 63.4 89.9 9.7 54.3 (+0.8)
+Averaged CLS attn. 63.4 90.2 9.9 54.5 (+1.0)
+Patch saliency [6] 62.9 90.3 9.6 54.2 (+0.7)
+ATS [9] 63.5 90.3 9.8 54.5 (+1.0)
+(b) Context token summarization strategy.
+Case HMDB UCF SSv2 All ($\Delta$)
+Baseline 62.6 89.2 8.7 53.5
+No merge 57.2 85.6 7.7 50.2 ($-3.3$)
+Random merge 58.8 87.1 7.5 51.2 ($-2.3$)
+K-means [26] 62.1 89.7 9.0 53.6 (+0.1)
+DPC-KNN [14] 63.3 90.2 9.8 54.4 (+0.9)
+Bipartite soft matching [16, 1] 63.4 90.2 9.9 54.5 (+1.0)
+Bipartite w/ attention weights 62.9 89.8 9.9 54.2 (+0.7)
+Bipartite w/ saliency weights [6] 62.4 89.9 9.6 54.0 (+0.5)
+Table 9: TC design ablation. We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. Bias is defined in Eq. (8).
+(a) Positional embedding design.
 Case HMDB UCF SSv2 All
-No selection 62.8 89.8 9.7 54.1
-Head-wise key norm 62.3 89.8 9.8 54.0
-Averaged key norm 62.5 89.4 9.3 53.7
-Head-wise CLS attention 63.4 89.9 9.7 54.3
-Averaged CLS attention 63.4 90.2 9.9 54.5
-(b) Token aggregation method.
-Case HMDB UCF SSv2 All
-No merge 57.2 85.6 7.7 50.2
-Temporal tube 57.9 86.4 8.3 50.9
-Random merge 58.8 87.1 7.5 51.2
-K-means 61.7 89.7 9.0 53.4
-Bipartite matching 63.4 90.2 9.9 54.5
-(c) Seed token ratio $\alpha$.
+Spatial embedding 62.9 90.0 9.8 54.2
+Joint space-time embedding 63.2 90.2 9.8 54.4
+Spatial embedding + Bias 63.4 90.2 9.9 54.5
+Joint embedding + Bias 62.9 90.2 9.8 54.3
+(b) Seed token ratio $\alpha$.
 $\alpha$ HMDB UCF SSv2 All
@@
 0.6 63.1 90.2 9.8 54.4
-(d) Context token $k$.
+(c) Context token $k$.
 $k$ HMDB UCF SSv2 All
@@
 128 62.8 90.1 9.9 54.3
-256 60.9 89.5 9.3 53.2
-(e) Positional embedding.
-Case HMDB UCF SSv2 All
-SE 62.9 90.0 9.8 54.2
-JE 63.2 90.2 9.8 54.4
-SE+Bias 63.4 90.2 9.9 54.5
-JE+Bias 62.9 90.2 9.8 54.3
-5 Analysis and Discussion
-This section examines the design choices and impact of each component in our model: Temporal Contextualization (TC) and Video-conditional Prompting (VP).
-We mainly adopt the zero- and few-shot settings and report the average of top-1 accuracy from $K=2,4,8,16$ for the $K$-shot setup.
-5.1 Component-wise Ablation
-Table 5 shows the impact of TC, VP, and weight-space ensembling (WE) on our baseline in the zero-shot setting.
-(a) Adopting TC with WE leads to an average gain of 3.3%p, while (b) incorporating VP with WE results in a 2.7%p improvement, highlighting the significance of TC in our framework.
-(c) When both VP and TC are applied to the baseline, we observed a 1.6%p improvement on average.
-Finally, by integrating all three components, TC-CLIP achieves a total 3.7%p gain on average.
-5.2 Temporal Contextualization (TC) Ablation
-Table 6 analyzes the design choice of TC. Experiments are conducted on the few-shot setting using the baseline model with TC.
-Seed token selection metric.
-Using [CLS] attention score averaged from multiple heads is the best choice, resulting in a 0.4%p gain on average compared to the no-selection case, i.e., using all tokens for aggregation.
-Token aggregation methods.
-Using seed tokens without merging leads to inferior performance, encountering the extrapolation issue.
-When comparing naive aggregation methods, random merge surpasses temporal tube since random merge has more chance to aggregate relevant tokens while temporal tube are enforced to group tokens of the same spatial positions.
-Finally, we compared the clustering algorithms and selected bipartite matching as our default choice.
-Seed token selection ratio $\alpha$. While TC is not sensitive to the choice of $\alpha$, we picked $\alpha=0.3$ as our default value, i.e., using 30% of total tokens.
-Context token number $k$.
-By varying the value of $k$, we observed that selecting a modest amount of merging degree is important for achieving the best performance.
-An excessively low $k$ leads to the lost of spatial details like part-wise information, while an overly high $k$ results in inadequate contextualization of tokens into relevant sets.
-Therefore, we choose $k=96$ as our default setting.
-Positional embedding.
-Among spatial embedding (SE), joint space-time embedding (JE), and our proposed learnable bias (Eq.
-10), we observe that using SE with bias yields the best result.
-We conjecture that the bias effectively consolidates the local frame-level information and global video-level information in a layer-wise and head-wise manner.
-Table 7: Video-conditional Prompting (VP) ablation. Default settings are marked in gray.
-(a) Vision token selection (few-shot).
-Case HMDB UCF SSv2 All
-CLS tokens 63.3 90.3 9.8 54.4
-GAP tokens 63.7 90.5 9.8 54.7
-Context tokens 63.9 90.7 9.8 54.8
-(b) Number of prompt vectors (few-shot).
-Case HMDB UCF SSv2 All
-2 62.0 90.1 9.4 53.8
-4 63.9 90.7 9.8 54.8
-8 63.7 90.4 9.8 54.7
-(c) Text prompting design (zero-shot). All the models are evaluated without weight ensembling.
+Table 10: Text prompting design ablation on the zero-shot setting. All the models are evaluated without the weight ensemble.
 Case Use context tokens? HMDB-51 UCF-101 K-600 All ($\Delta$)
@@
 (d) Vision-text late-fusion 53.7 $\pm$ 0.7 79.0 $\pm$ 0.7 70.9 $\pm$ 0.6 67.9 (+0.6)
-5.3 Video-conditional Prompting (VP) Ablation
-In Table 7, we analyze the design choice of the VP module in TC-CLIP on the few-shot and zero-shot setting.
-Vision token selection.
-We compare using the context tokens with using [CLS] or Global Average Pooled (GAP) tokens from every frame.
-Using context tokens yields the best result, demonstrating that the proper contextualization from the video is essential for transferring the video information to the text side.
-Number of prompt vectors. Increasing the number of prompt vectors does not necessarily improve performance. 4 prompt vectors are employed by default.
+3.2 Analysis and Discussion
+This section examines the design choices and impact of each component in our model: Temporal Contextualization (TC) and Video-conditional Prompting (VP).
+We mainly adopt the zero- and few-shot settings and report the average of top-1 accuracy with $K=2,4,8,16$ for the $K$-shot setup.
+In addition to the analyses discussed in this subsection, we present more analyses and qualitative results in the supplementary document.
+Component-wise ablation.
+Table 7 shows the impact of TC and VP on our baseline in the zero-shot setting.
+Integrating TC gives an average gain of 0.7%p over the baseline and the gap increases to 1.6%p after adopting WE; WE is more favorable to our approach than the baseline.
+Adopting VP also leads to a substantial gain of 1.1%p, highlighting its own contribution.
+When both VP and TC are applied to the baseline, an average improvement goes up to 1.6%p, which finally leads to 2.0%p gain after applying WE.
+Token aggregation strategies.
+In Table 8, we verify the effectiveness of TC across diverse token aggregation methods.
+Experiments are conducted on the few-shot setting using the baseline model with TC.
+(a) While TC still works well without token selection, we observe that collecting informative seed tokens based on token importance, such as attention or saliency scores, improves the quality of encoded tokens by suppressing the background.
+(b) Directly using the seed tokens without merging reduces performance due to the extrapolation issue.
+The degradation with random merging also highlights the requirement of token clustering based on relevance.
+Finally, consistent gains from various token merging approaches verify the robustness of TC regardless of algorithms.
+Positional embedding.
+Table 9(a) shows that using the proposed learnable bias (Eq.
+(8)) with spatial positional embedding yields the best result.
+We conjecture that the bias effectively consolidates the local frame-level information and global video-level information in a layer-wise and head-wise manner.
+Number of seed and context tokens.
+While TC is not sensitive to the choice of $\alpha$, as shown in Table 9(b), we picked $\alpha=0.3$ as our default value, i.e., using 30% of total tokens as seed tokens.
+In Table 9(c), the context token number $k$ is chosen to set a modest amount of merging degree.
 Text prompting design.
-We observe that (a) naively adding learnable prompt vectors without video instance conditioning does not enhance the zero-shot transferability and decreases the average accuracy by 0.2%p.
-(b) In contrast, employing VP design with [CLS] tokens consistently improves the accuracy across all datasets, and (c) using context tokens further enhances the performance, resulting in a 1.6%p gain.
+In Table 10, we observe that (a) a naïve integration of learnable prompt vectors without video instance conditioning is not particularly helpful for the zero-shot transferability, rather decreasing the average accuracy.
+In contrast, (b) employing VP design with [CLS] tokens consistently improves the accuracy across all datasets, and (c) using context tokens further enhances the performance, resulting in a 1.6%p gain.
 We also compare VP with (d) vision-text late-fusion design, i.e., the cross-attention of context tokens and the final representation of the text embedding.
-This design shows worse performance in UCF-101 and K-600 datasets than our VP, demonstrating the superiority of our design choice.
-5.4 Qualitative Analysis
+This design performs worse in UCF-101 and K-600 datasets than our VP, verifying the effectiveness of our design choice.
+Figure 6: Context token visualization. TC-CLIP selects the informative seed tokens and summarizes them into context tokens across frames. The disc (red) is merged into one token over the video.
+Figure 7: Attention visualization. While ViFi-CLIP fails to attend to the hands moving away and misinterprets the action as colliding, TC-CLIP correctly predicts by exploiting temporal consistency.
 Context token visualization.
@@
 6 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP.
-In this video, the informative regions regarding the action of “disc golfing” in each frame, including the person and the disc, are selected as seed tokens.
-To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region.
+In this video, the informative regions regarding the action of disc golfing in each frame, including the person and the disc, are selected as seed tokens.
+To visualize each context token, we colorize its corresponding source token positions using the average color of the input video patches of that region.
 Note that a single context token (highlighted in red) successfully tracks the disc across multiple frames.
@@
 Fig.
-7 visualizes the attention map of ViFi-CLIP [26] and our TC-CLIP on the SSv2 dataset.
-In this video where two hands grab objects and then move away, ViFi-CLIP [26] fails to attend to the hands from the middle of the sequence and misinterprets the action as “colliding with each other”.
-In contrast, our TC-CLIP considers the temporal context across the sequence by its design, and thus consistently attends to the hands throughout the entire video and correctly predicts the action as “moving away from each other”.
-See more visualizations in Appendix 0.E.
-Figure 6: Context token visualization. TC-CLIP selects the informative seed tokens and summarizes them into context tokens across frames. The disc (red) is merged into one token over the video.
-Figure 7: Attention visualization.
-While ViFi-CLIP fails to attend to the movement of hands moving farther away and misinterprets the action as colliding, TC-CLIP correctly predicts with temporal consistency.
+7 visualizes the attention map of ViFi-CLIP [35] and TC-CLIP on the SSv2 dataset.
+In this video, where two hands grab objects and then move away, ViFi-CLIP [35] fails to attend to the hands from the middle of the sequence and misinterprets the action as colliding with each other.
+In contrast, TC-CLIP considers the temporal context across the sequence by its design, and thus consistently attends to the hands throughout the entire video and correctly predicts the action as moving away from each other.
+4 Related Work
+Token aggregation.
+Recent studies on token aggregation aim to reduce the number of tokens given to image Transformers [34, 21, 31, 24, 48, 49, 9, 27, 29, 51, 1, 20] and video Transformers [37, 40, 36, 23, 7, 6] for efficient inference.
+While some of these approaches train additional networks for token selection [34, 21, 40] or fusion [31, 37], we focus on parameter-free approaches, categorized into token pruning and merging.
+Pruning-based methods [24, 48, 49, 9] eliminate uninformative tokens by measuring their importance using a metric such as a self-attention score, whereas merging-based methods combine tokens with large semantic similarity into single units using clustering algorithms such as $k$-means [29], DPC-KNN [51], and bipartite soft matching [1, 36, 23].
+To minimize information loss, several studies [27, 7, 6] consider both token importance and similarities as aggregation criteria.
+While our primary goal is not to improve efficiency, we employ both pruning and merging techniques to connect relevant tokens and summarize essential contexts within videos.
+Although there are semantic segmentation studies [46, 47, 28] that link relevant spatial data, they rely on learnable cluster centers with slot- or cross-attention blocks, and thus differ from our approach.
+Prompt learning.
+Several studies on prompt learning [13, 53, 52, 18, 35, 43, 15] transfer VLMs to downstream tasks by optimizing a discrete set of prompt vectors.
+In video recognition, [15] has introduced text prompt tuning, while ViFi-CLIP [35] and Vita-CLIP [43] perform prompting in both vision and text branches.
+However, these prompt vectors are separately optimized and not shared across the modalities.
+In image recognition, Co-CoOp [52] performs an instance-conditional prompt tuning by explicitly conditioning text prompts on the [CLS]
+tokens from image instances.
+MaPLe [18] learns multi-modal prompting by sharing layerwise context prompts for both branches.
+Unlikely, we generate video-conditional prompts by utilizing contextualized tokens as vision inputs and injecting summarized video information into text prompt vectors.
 5 Conclusion
 We have introduced TC-CLIP, a novel video understanding paradigm that leverages holistic video information within the encoding process.
-Unlike prior approaches that access only a limited range of tokens, our temporal contextualization summarizes informative tokens from the entire video and utilizes them during the attention operation.
-While these tokens are exploited to infuse temporal information on the vision side, they also serve as a source for video-conditional text prompting, enhancing instance-wise context on the text side.
-Extensive experiments and analysis on diverse video recognition benchmarks and evaluation protocols demonstrate the superiority of TC-CLIP and its design choices.
-Limitation. Due to resource constraints, we verified TC-CLIP only using ViT-B. We will further explore its scalability with ViT-L and ViT-H.
+Unlike prior approaches that access only a limited range of tokens, the proposed temporal contextualization summarizes informative tokens from the entire video and utilizes them for attention operations.
+While these tokens are employed to infuse temporal information on the vision side, they also serve as a source for video-conditional text prompts, thus enhancing the instance-wise context on the text side.
+Extensive experiments and analyses on diverse benchmarks and evaluation protocols demonstrate the superiority of TC-CLIP and justify its design choices.
+Acknowledgements
+Experiments are based on the NAVER Smart Machine Learning NSML [19] platform.
+This research was partly supported by the Bio & Medical Technology Development Program of the National Research Foundation (NRF) (No.
+2021M3A9E4080782) and the IITP grants [No.
+RS-2021-II212068; No.
+RS-2021-II211343] funded by the Korean government (MSIT).
 Appendix 0.A Fine-tuning with the Kinetics-400 Pretrained Model
@@
 TC-CLIP (Ours) 79.4 58.3 67.2 97.5 84.5 90.5 19.6 15.6 17.4 65.5 52.8 58.5
-We additionally compare TC-CLIP with the state-of-the-art methods on few-shot and base-to-novel settings using the Kinetics-400 pretrained model.
-In this setting, all the models are first pretrained on the Kinetics-400 dataset and subsequently fine-tuned on each dataset.
-Table 8 summarizes the comparison results in the few-shot action recognition regime.
-Notably, TC-CLIP surpasses other methods by a large margin of at least 2%p in the average.
-Furthermore, Table 9 shows the comparison results in the base-to-novel generalization setup.
-In an average over all datasets, TC-CLIP shows distinct gaps from the baseline model ViFi-CLIP [26] by 3.5%p, 6%p, and 4.8%p in base, novel, and harmonic mean (HM), respectively.
-Table 10: Temporal subset analysis using the temporal subset [27] on Kinetics-400 and SSv2. Gains over ViFi-CLIP are indicated in green.
-K-400 fully-supervised SSv2 16-shot
-Method All Temporal All Temporal
-ViFi-CLIP [26] 83.9 87.8 12.4 25.9
-TC-CLIP (Ours) 85.2 ($+$1.3) 89.2 ($+$1.4) 14.0 ($+$1.6) 29.9 ($+$4.0)
-Table 11: VP layer selection ablation.
-We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. The late-stage prompting design generalizes well.
+Tables 11 and 12 present the comparison results using the K-400 pretrained model on the few-shot and base-to-novel settings.
+All the models are first pretrained on the Kinetics-400 dataset and subsequently fine-tuned on each dataset.
+TC-CLIP demonstrates superior performance over all other methods by significant margins.
+Particularly in the base-to-novel setup, TC-CLIP outperforms ViFi-CLIP [35] with notable gaps of 3.5%p, 6%p, and 4.8%p in the base, novel, and harmonic mean (HM) on average, respectively.
+Appendix 0.B More Ablation Study on VP
+Table 13: Video-conditional Prompting (VP) ablation. We report $K$-shot training results where the top-1 accuracy in each dataset is averaged over $K=2,4,8,16$. Default settings are marked in gray.
+(a) Number of prompt vectors.
+Case HMDB UCF SSv2 All
+2 62.0 90.1 9.4 53.8
+4 63.9 90.7 9.8 54.8
+8 63.7 90.4 9.8 54.7
+(b) Vision input token selection.
+Case HMDB UCF SSv2 All
+CLS tokens from all frames 63.3 90.3 9.8 54.4
+GAP tokens from all frames 63.7 90.5 9.8 54.7
+Context tokens 63.9 90.7 9.8 54.8
+(c) Input layer selection.
 $L_{\text{text}}$ $L_{\text{vision}}$ HMDB UCF SSv2 All
@@
 12 12 63.9 90.7 9.8 54.8
-Table 12: VP initialization ablation by varying initialization cases of the VP layer and prompt vectors.
-We report $K$-shot training results where the top-1 accuracy is averaged over $K=2,4,8,16$.
-In general, it is advantageous to avoid random initialization.
+(d) Layer and prompt initialization.
 Layer init. Prompt init. HMDB UCF SSv2 All
-CLIP a photo of a 63.9 90.7 9.8 54.8
-Random a photo of a 63.5 90.5 9.9 54.6
-CLIP Random 62.5 90.2 9.5 54.1
-Appendix 0.B Temporal Subset Analysis
+CLIP weight a photo of a 63.9 90.7 9.8 54.8
+Random init. a photo of a 63.5 90.5 9.9 54.6
+CLIP weight Random init. 62.5 90.2 9.5 54.1
+Table 14: Computational cost–performance trade-off of VP design.
+In the case of vision-text late-fusion design, class name embeddings are pre-computed.
+Models are evaluated on the zero-shot setting without the weight ensemble.
+Costs are measured using a single A6000 GPU.
+K-600 GFLOPs Latency (s)
+Case Top-1 Vision Text Cross-modal Training Inference
+Vision-text late-fusion 70.9 301 2.96 0.06 0.54 (1.00$\times$) 0.042 (1.00$\times$)
+Video-conditional Prompting (VP) 72.7 (+1.8) 301 2.96 0.06 0.58 (1.07$\times$) 0.045 (1.07$\times$)
+Table 13 examines the design choice of VP in TC-CLIP on the few-shot setting.
+Number of prompt vectors. Increasing the number of prompt vectors does not necessarily improve performance. 4 prompt vectors are employed by default.
+Vision token selection.
+Using context tokens in VP yields better results than employing [CLS] tokens or global average pooled (GAP) tokens from all frames.
+This demonstrates that proper contextualization of vision features is essential to transfer the video information to the text side.
+Input layer selection.
+We vary the layer indices of the text and vision inputs $\{L_{\text{text}},L_{\text{vision}}\}$ in the VP module $f_{\theta_{\text{VP}}}(\mathbf{p}^{L_{\text{text}}-1},\mathbf{s}^{L_{\text{vision}}}_{\text{proj}})$.
+We observe that conditional prompting at the early stage ($L_{\text{text}}=1$) does not generalize well, regardless of the vision layer index.
+The early-stage prompting design is hard to generalize in a full fine-tuning scenario, possibly because CLIP was initially trained in a vision-text late-alignment fashion.
+Consequently, we choose the late-stage prompting by adopting the last layers for both modalities.
+Layer and prompt initialization.
+We initialize the VP module’s weight using the weight from the last layer of the CLIP text encoder because random initialization often results in unstable training results in the few-shot scenario.
+Similarly, it is beneficial to initialize the learnable prompt vectors using the prompt template “a photo of a” following several prompt tuning methods [52, 18].
+Computational cost analysis.
+Although VP requires the instance-conditional computation of text embeddings, the added cost is minor.
+As in Table 14, for a pair of video and text inputs, the GFLOPs required by the text encoder cost only about 1% of those needed by the vision encoder.
+Given these minimal text-related costs, VP adds only an extra 0.07$\times$ in latency compared to the vision-text late-fusion design using pre-computed text embeddings.
+Considering the observed performance gain, this is an acceptable trade-off.
+Table 15: Comparison with state-of-the-arts on zero-shot action recognition using ViT-L/14.
+All the models are trained on Kinetics-400 and directly evaluated on other datasets.
+† denotes that the results are reproduced with our implementation.
+The best results are in bold-faced numbers, and the second-best ones are underlined.
+Method WE HMDB-51 UCF-101 K600 (Top-1) K600 (Top-5) All (Top-1)
+ViFi-CLIP [35] 55.6 $\pm$ 0.5 86.1 $\pm$ 0.8 77.8 $\pm$ 0.9 95.6 $\pm$ 0.2 73.2
+TC-CLIP (Ours) 56.1 $\pm$ 0.3 86.9 $\pm$ 0.9 80.1 $\pm$ 0.7 96.5 $\pm$ 0.1 74.4
+ViFi-CLIP [35] 55.8 $\pm$ 0.7 88.1 $\pm$ 1.3 81.1 $\pm$ 0.7 96.7 $\pm$ 0.1 75.0
+Open-VCLIP [44] 59.0 $\pm$ 0.6 87.6 $\pm$ 1.2 81.1 $\pm$ 0.8 96.3 $\pm$ 0.3 75.9
+TC-CLIP (Ours) 57.1 $\pm$ 0.7 88.9 $\pm$ 0.9 83.1 $\pm$ 0.7 97.3 $\pm$ 0.1 76.4
+Table 16: Temporal subset analysis using the temporal subset [38] on Kinetics-400 and SSv2. Gains over ViFi-CLIP are indicated in green.
+K-400 fully-supervised SSv2 16-shot
+Method All Temporal All Temporal
+ViFi-CLIP [35] 83.9 87.8 12.4 25.9
+TC-CLIP (Ours) 85.2 ($+$1.3) 89.2 ($+$1.4) 14.0 ($+$1.6) 29.9 ($+$4.0)
+Appendix 0.C Scalability with ViT-L/14
+Table 15 shows the zero-shot performance comparison using CLIP ViT-L/14 as a backbone.
+In the case of using WE, our model outperforms ViFi-CLIP [35] and Open-VCLIP [44] by 1.4%p and 0.5%p on average, respectively.
+Appendix 0.D Temporal Subset Analysis
 We adopt the temporal subset analysis suggested in [38] to further analyze the temporal modeling ability of trained models.
-In this setting, the temporal subset consists of specific action classes that require more temporal information to recognize them.
-After shuffling frames in time and measuring whether the human annotator can still recognize the action, the classes that cannot be recognized are added to the temporal subset.
-The resulting temporal classes in Kinetics-400 and SSv2 are a total of 32 classes and 16 classes, respectively.
-Table 10 compares the top-1 accuracy using full validation splits and the temporal subsets.
-We use fully-supervised models for Kinetics-400 and 16-shot-trained models for SSv2.
-Our model shows unique gains over ViFi-CLIP [26] for the temporally challenging videos, more than when evaluated across the entire validation set.
-This demonstrates the superiority of TC-CLIP in handling temporal information.
-Appendix 0.C More Ablation Study on VP
-VP layer selection.
-Table 0.A presents the ablation study on the layer selection in the VP module.
-We compare the few-shot classification accuracy by varying the layer indices of the text and vision inputs $\{L_{\text{text}},L_{\text{vision}}\}$ in the VP module $f_{\theta_{\text{VP}}}(\mathbf{p}^{L_{\text{text}}-1},\mathbf{s}^{L_{\text{vision}}}_{\text{proj}})$.
-We observe that applying conditional prompting at the early stage ($L_{\text{text}}=1$) does not generalize well, regardless of the vision layer index.
-This suggests the early-stage prompting design is hard to generalize in a full fine-tuning scenario, possibly because CLIP was initially trained in a vision-text late-alignment fashion.
-The middle- and late-stage prompting designs are also compared, and the last layer indices are finally chosen for both modalities.
-VP initialization.
-Table 0.A presents the ablation study on the initialization strategy of the VP layer and learnable prompt vectors.
-By default, we initialize the weight of the VP module using the weight from the last layer of the CLIP text encoder since randomly initializing it occasionally incurs unstable training results in the few-shot scenario.
-Similarly, it is beneficial to initialize the learnable prompt vectors using the prompt template “a photo of a” as generally done in the prompt tuning methods [37, 17].
-Appendix 0.D Impact of Weight-space Ensemble
+The temporal subset consists of several action classes that require more temporal information to recognize them, i.e., the classes of videos that cannot be recognized by human annotators after randomly shuffling the frames.
+As shown in Table 16, TC-CLIP’s gains over ViFi-CLIP [35] on the temporal subsets are more substantial than the gains when evaluated on the full validation splits, demonstrating the superiority in handling temporal information.
+Appendix 0.E Impact of Weight-space Ensembling
 Figure 8: Weight averaging ablation.
@@
 Specifically, the backbone weights of both vision and text encoders are linearly interpolated between CLIP and fine-tuned model, i.e., $\theta_{w}=(1-w)\cdot\theta_{\text{CLIP}}+w\cdot\theta_{\text{fine-tuned}}$.
-The $y$-axis displays the average accuracy on the zero-shot video datasets and the $x$-axis displays the accuracy on the fine-tuning dataset K-400.
-Our TC-CLIP achieves a better trade-off than the baseline as our curve is always on top of the baseline’s curve.
+The $y$-axis shows the average accuracy on the zero-shot video datasets, and the $x$-axis means the accuracy on the fine-tuning dataset K-400.
+Our model achieves a better trade-off than the baseline as our curve is always on top of the baseline’s curve.
 This demonstrates that our model takes more advantages from weight ensembling.
@@
 Regions highlighted in red represent a specific object or part grouped into a single context token throughout the video.
-Figure 10: Attention visualization of TC-CLIP in comparison with ViFi-CLIP [26] on Kinetics-400, Kinetics-600, and SSv2 datasets.
+Figure 10: Attention visualization of TC-CLIP in comparison with ViFi-CLIP [35] on Kinetics-400, Kinetics-600, and SSv2 datasets using [CLS] token as a query in each frame.
 (a)–(b): TC-CLIP tends to focus more on fast-moving parts such as hands and arms.
@@
 (e)–(f): TC-CLIP consistently attends to the main object with deformations throughout the video.
-Appendix 0.E More Visualizations of Context Tokens and Attentions
+Figure 11: Attention visualization of TC-CLIP in comparison with various temporal information learning approaches on SSv2 dataset.
+We visualize the attention map in the last vision encoder layer using a ball (top) and a hand (bottom) as a query (denoted with red boxes).
+To visualize the attention map from TC, we assign attention values of context tokens to their corresponding source patch token positions.
+Unlike other approaches, our method successfully highlights informative regions globally over frames.
+Appendix 0.F More Visualizations of Context Tokens and Attentions
 Context token visualization.
@@
 9 visualizes the seed tokens and context tokens from the last layer of the vision encoder in TC-CLIP.
-As can be seen, the seed tokens mainly consist of patch tokens from the most informative regions in each frame, which contain the foreground, such as a person, animals, hands, and objects.
+The seed tokens mainly consist of patch tokens from the most informative regions in each frame, often corresponding to the foreground, such as a person, animals, hands, and objects.
 To visualize each context token, we colorize its corresponding source token positions using the average color of the input image patches of that region.
 It is noteworthy that a single context token (highlighted in red) successfully tracks and summarizes a specific object or part throughout the entire video.
-Attention visualization.
+Class token attention visualization.
 Fig.
-10 visualizes the attention maps of TC-CLIP compared to ViFi-CLIP [26].
+10 visualizes the attention maps of TC-CLIP compared to ViFi-CLIP [35] using [CLS] token as a query in each frame.
 As shown in Fig.
-10(a)–(b), TC-CLIP tends to focus more on dynamically moving parts, such as hands and arms, during the action of throwing or shooting objects.
-Furthermore, TC-CLIP attends to multiple objects simultaneously based on inter-object relationships, as illustrated in Fig.
-10(c)–(d).
-For example, during actions like “swinging baseball bat,” it effectively attends to both the bat and the baseball being struck.
-This highlights the TC-CLIP’s capability to understand complex interactions based on a temporal context in dynamic scenes, which is potentially more helpful than ViFi-CLIP, which only attends to the most salient region in each frame.
-Finally, Fig.
-10(e)–(f) shows TC-CLIP’s consistent attention towards objects with deformations across video frames, which are more striking than ViFi-CLIP’s.
-Appendix 0.F Experimental Setup Details
-0.F.1 Dataset Details
+10(a)–(b), during the action of throwing or shooting objects, TC-CLIP tends to focus more on dynamically moving parts such as hands and arms.
+Furthermore, as in Fig.
+10(c)–(d), TC-CLIP highlights multiple objects simultaneously based on inter-object relationships.
+During actions like “swinging baseball bat,” TC-CLIP focuses on both the bat and the baseball being struck, whereas ViFi-CLIP only highlights salient areas in individual frames.
+Fig.
+10(e)–(f) also shows TC-CLIP’s consistent attention towards objects with deformations across frames, which is more striking than ViFi-CLIP’s.
+Patch token attention visualization.
+Fig.
+11 shows the attention maps of TC-CLIP compared to other temporal modeling approaches [30, 43, 44] by using a patch token as a query.
+To visualize the attention map from TC, we assign attention values of context tokens to their corresponding source patch token positions.
+In both examples, the token interactions of cross-frame attention [30, 43] and temporal window expansion [44] cannot reach the frames far from the query position, although the main action actually happens in the latter part of videos.
+The joint space-time attention model, on the other hand, is capable of global modeling but fails to focus on informative regions.
+In contrast, TC-CLIP consistently highlights the regions relevant to the query positions (e.g., hands and grabbed objects) throughout the video, leading to more accurate predictions.
+Appendix 0.G Experimental Setup Details
+0.G.1 Dataset Details
 We conduct experiments over 5 action recognition benchmarks: Kinetics-400 [17] & 600 [2], HMDB-51 [22], UCF-101 [39], and Something-Something v2 (SSv2) [10].
@@
 The standard split consists of 168,913 training videos and 24,777 validation videos.
-0.F.2 Implementation Details
-During the bipartite matching, we start with the seed tokens arranged based on the [CLS] token attention values in each frame.
+0.G.2 Implementation Details
+During the bipartite soft matching [16, 1], we start with the seed tokens arranged based on the [CLS] token attention values in each frame.
 These tokens are then divided into two sets by alternating positions.
-Subsequently, $r$ pairs of tokens with the highest cosine similarity are merged by averaging their features, and remaining two sets are then concatenated back together.
+Subsequently, $r$ pairs of tokens with the highest cosine similarity are merged by averaging their features, and the remaining two sets are then concatenated back together.
+We set $r$ to 100 in practice.
 This process is repeated iteratively, employing a constant $r$ scheduling for every iteration with an exception in the final iteration to ensure that the number of final context tokens becomes $k$.
-Following [37, 17], the learnable prompts are initialized with the prompt “a photo of a”.
+During the training, we sample 16 frames to form a video clip.
+During the evaluation, two temporal clips with one spatial crop (2 $\times$ 1 view) per video are sampled to produce a prediction unless otherwise stated.
+The learnable prompts are initialized with the prompt “a photo of a” following [52, 18], and the weight of the VP module is initialized with the weight from the last layer of the CLIP text encoder.
 For training recipes, we follow [35] for zero-shot, few-shot, and fully-supervised settings and follow [11] for base-to-novel generalization.
 By default, we use the AdamW optimizer with momentum betas of (0.9, 0.98) and a weight decay of 0.001.
-The initial learning rate of the VP module is 10×10$\times$10 × larger than the base learning rate in each setting.
+The VP module’s initial learning rate is 10×10$\times$10 × larger than the base learning rate in each setting.
 Training configurations and evaluation metrics in each protocol are specified below.
@@
 Few-shot action recognition.
-We adopt the $K$-shot training splits from [26] that randomly samples $K=2,4,8,16$ videos from each class on HMDB-51, UCF-101, and SSv2.
+We adopt the $K$-shot training splits from [35] that randomly sampled $K=2,4,8,16$ videos from each class on HMDB-51, UCF-101, and SSv2.
 The models are evaluated using the first validation split of HMDB-51 and UCF-101 and the full validation split of SSv2.